### v1

# Inverse Reinforcement Learning by Estimating Expertise of Demonstrators

Abstract
In Imitation Learning (IL), utilizing suboptimal and heterogeneous demonstrations presents a substantial challenge due to the varied nature of real-world data. However, standard IL algorithms consider these datasets as homogeneous, thereby inheriting the deficiencies of suboptimal demonstrators. Previous approaches to this issue typically rely on impractical assumptions like high-quality data subsets, confidence rankings, or explicit environmental knowledge. This paper introduces IRLEED, Inverse Reinforcement Learning by Estimating Expertise of Demonstrators, a novel framework that overcomes these hurdles without prior knowledge of demonstrator expertise. IRLEED enhances existing Inverse Reinforcement Learning (IRL) algorithms by combining a general model for demonstrator suboptimality to address reward bias and action variance, with a Maximum Entropy IRL framework to efficiently derive the optimal policy from diverse, suboptimal demonstrations. Experiments in both online and offline IL settings, with simulated and human-generated data, demonstrate IRLEED’s adaptability and effectiveness, making it a versatile solution for learning from suboptimal demonstrations.
Learning from Suboptimal Demonstrations, Inverse Reinforcement Learning
Figure 1: IRLEED can be applied in the suboptimal demonstration setting to estimate the ground truth reward $r$, which is used to find the optimal policy. Left: A heterogeneous dataset is collected from multiple sources with varying degrees of optimality. We categorize this optimality by using accuracy to represent the reward bias, and precision to represent the variance in action choices. Right: We can infer the demonstrator policies using a model for suboptimal behaviour based on the Boltzmann rationality principle, which captures both the accuracy $\epsilon_{i}$, and the precision $\beta_{i}$, as compared to the ground truth reward $r$. Middle: Using estimates of the demonstrator policies $\hat{\pi}_{i}$ along with the demonstrations $\mathcal{D}_{i}$, we can optimize for the true reward $r$, as well as the parameters that capture accuracy $\epsilon$ and precision $\beta$.
1 Introduction
Reinforcement Learning (RL) has proven to be a versatile and powerful tool across a wide range of applications, from navigating autonomous vehicles, to playing complex games like Go and Chess. The efficacy of RL algorithms is largely attributed to their ability to optimize hand-crafted reward functions (Brockman et al., 2016). However, designing these functions is often challenging and impractical in complex environments (Hadfield-Menell et al., 2017). A prevalent alternative to this dilemma is Imitation Learning (IL), which bypasses the need for cumbersome reward engineering by leveraging expert demonstrations to instill desired behaviors (Argall et al., 2009).
The two main approaches utilized in IL are: behavioral cloning (Pomerleau, 1991), which acquires a policy through a supervised learning approach using state-action pairs provided by the expert, and Inverse Reinforcement Learning (IRL) (Ng & Russell, 2000), which solves for the reward function that makes the expert’s behavior optimal, subsequently facilitating the training of an IL policy. Despite the effectiveness of behavior cloning in simple environments with large amounts of data, learning a policy to fit single time step decisions leads to compounding errors due to covariance shift (Ross & Bagnell, 2010). On the other hand, IRL learns a reward function that prioritizes entire trajectories over others, considering the sequential nature of the decision making problem at hand (Abbeel & Ng, 2004). Consequently, the success of IRL has spurred the development of IL techniques that either explicitly or implicitly incorporate environment dynamics (Ho & Ermon, 2016; Fu et al., 2018; Kostrikov et al., 2019; Garg et al., 2021).
However, a critical assumption in these methodologies is the availability of high-quality demonstration data. In many practical situations, especially with crowd-sourced or varied data, the quality of demonstrations is inconsistent (Mandlekar et al., 2021; Belkhale et al., 2023). For example, foundation models like GPT-3 which are trained on large and uncurated datasets, exhibit limitations such as generating biased outputs that negatively stereotype certain groups (Abid et al., 2021). In robotics, data curation becomes vital, since utilizing smaller real world datasets with noise or biases can lead to hazardous situations. Furthermore, assuming uniformity in demonstration quality overlooks the unique intentions of individual demonstrators, potentially leading to suboptimal learning outcomes (Eysenbach et al., 2018). Therefore, it is critical to develop IL methods that can account for the heterogeneity and suboptimality of demonstration data.
Addressing the challenges posed by learning from suboptimal and heterogeneous demonstrations is complex. One line of research focuses on heterogeneous demonstrations without considering their quality (Chen et al., 2020, 2023), assuming that all provided demonstrations are optimal. Other works consider suboptimal demonstrations, but require unrealistic prerequisites such as explicit knowledge of environment dynamics (Brown et al., 2020; Chen et al., 2021), a set of confidence scores over the demonstrations (Zhang et al., 2021; Cao & Sadigh, 2021; Wu et al., 2019; Kuhar et al., 2023), or direct access to a subset of expert demonstrations (Xu et al., 2022; Yang et al., 2021). A recently proposed approach that avoids these assumptions is ILEED (Beliaev et al., 2022), which directly models the expertise of demonstrators within a supervised learning framework, thus enabling one to learn from the optimal behavior while filtering out the suboptimal behavior of each demonstrator. While this method was proven effective, it is based on a behavioral cloning formulation, which overlooks the underlying environment dynamics, and its representation of suboptimal behavior is limited, presuming demonstrators to be noisy approximations of experts.
To overcome these challenges, this paper introduces IRLEED, Inverse Reinforcement Learning by Estimating Expertise of Demonstrators. As illustrated in Figure 1, IRLEED is a novel framework for IRL that accounts for demonstrator suboptimality without prior expertise knowledge. It comprises two core components: (1) a general model of demonstrator suboptimality based on the Boltzmann rationality principle, which captures both the reward bias and the variance in action choices of each demonstrator as compared to the optimal policy (2) a Maximum Entropy IRL framework, which is used to efficiently find the optimal policy when provided with a set of suboptimal and heterogeneous demonstrations. These elements enable IRLEED to effectively recover the ground truth policy from suboptimal and heterogeneous demonstrations, surpassing the limitations of previous models. Furthermore, IRLEED’s simplicity and compatibility with existing IRL techniques make it a versatile and powerful tool in the field of IL.
The main contributions of our work are:
  • We propose a novel framework designed to enhance existing IRL algorithms, addressing the challenges of learning from suboptimal and heterogeneous data.
  • We provide comparative insights against standard IRL and the behavior cloning approach, ILEED, showing how IRLEED generalizes both methods.
  • We empirically validate the success of our method against relevant baselines in both online and offline IL scenarios, utilizing both simulated and human datasets.
2 Preliminary
Notations We consider the standard Markov decision process (MDP) setting, defined by a tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{T},p_{0},r,\gamma)$, where $\mathcal{S}$, $\mathcal{A}$ represent state and action spaces, $\mathcal{T}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ represents the dynamics, $p_{0}:\mathcal{S}\rightarrow[0,1]$ represents the initial state distribution, $r\in\mathcal{R}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ represents the reward function, and $\gamma\in(0,1)$ represents the discount factor. In Section 3 and Section 4, we describe our method using finite state and action spaces, $\mathcal{S}$ and $\mathcal{A}$, but our experiments later utilize continuous environments. We use $\pi:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ to denote a policy that assigns probabilities to actions in $\mathcal{A}$ given states in $\mathcal{S}$. Considering the $\gamma$–discounted infinite time horizon setting, we define the expected value under a policy $\pi$ in terms of the trajectory it produces. Specifically, this expectation is represented as $\bar{f}_{\pi}=\operatorname{\mathbb{E}}_{\pi}[f(s,a)]$, which equals $\operatorname{\mathbb{E}}[\sum_{t=0}^{\infty}\gamma^{t}f(s_{t},a_{t})]$, where $s_{0}\sim p_{0},a_{t}\sim\pi(\cdot|s_{t})$, and $s_{t+1}\sim\mathcal{T}(\cdot|s_{t},a_{t})$. Additionally, we will use $\tilde{f}_{\mathcal{D}}=\operatorname{\mathbb{E}}_{\mathcal{D}}[f(s,a)]$ to denote the empirical expectation with respect to the trajectories $\tau_{j}=(s_{0},a_{0},\ldots,s_{T},a_{T})$ in demonstration set $\mathcal{D}=\{\tau_{j}\}_{j=1}^{M}$, where we leave out denoting trajectories of varying lengths for simplicity. Maximum Entropy IRL The goal of Inverse Reinforcement Learning (IRL) is to recover a reward $r\in\mathcal{R}$ that rationalizes the demonstrated behavior in dataset $\mathcal{D}$, where the provided dataset consists of trajectories $\tau$ sampled from an expert policy $\pi_{E}$. This general idea was first formulated as a feature-expectation matching problem (Abbeel & Ng, 2004): assuming that the feature vector $f:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}^{k}$ components quantify the expert’s behavior, to find a policy $\pi$ that performs equal to or better than the expert $\pi_{E}$, it suffices that their feature expectations match $\bar{f}_{\pi}=\bar{f}_{\pi_{E}}$. Unfortunately, this leads to an ill-posed problem as many policies can lead to the same feature counts. Maximum Entropy IRL resolves this ambiguity by choosing the policy which does not exhibit any additional preferences beyond matching feature expectations (Ziebart et al., 2008):
$\operatorname*{arg\,max}_{\pi}H(\pi),\textit{ such that: }\bar{f}_{\pi}=\tilde{f}_{\mathcal{D}},$ (1)
where in the infinite-time horizon setting, this equates to maximizing the discounted causal entropy $H(\pi)$ under the feature matching constraint (Bloem & Bambos, 2014).
In the MDP setting, the above problem is equivalent to finding the maximum likelihood estimate of $\theta$:
$\hat{\theta}=\operatorname*{arg\,max}_{\theta}\log\prod_{\tau_{j}\in\mathcal{D}}p_{0}(s_{0})\prod_{t=0}^{T}\pi_{\theta}(a_{t}|s_{t})\mathcal{T}(s_{t+1}|s_{t},a_{t}),$ (2)
with the parameterized policy $\pi_{\theta}$ defined recursively as:
[…]
The parameters $\theta\in\mathbb{R}^{k}$ specifying the soft Bellman policy, $\pi_{\theta}(a_{t}|s_{t})$, correspond to the dual variables of the feature matching constraint in Eq. (1). Without loss of generality, if we define the true reward signal as $r(s,a)=\theta^{\star\top}f(s,a)$ for some reward parameter ${\theta}^{\star}\in\mathbb{R}^{k}$ and features $f$, then the soft Bellman policy parameterized by $\theta^{\star}$ achieves the maximum possible return (for more details see (Ziebart, 2010), Theorem 6.2 and Corollary 6.11). [^1] To allow an arbitrary reward specification, the IRL objective defined in Eq. 1 can be expressed in terms of occupancy measures instead of feature expectations (Ho & Ermon, 2016). This formulates an iterative approach to finding the true reward parameter $\theta^{\star}$: given demonstration set $\mathcal{D}$ and your estimate $\theta$, alternate between inferring the policy $\pi_{\theta}(a_{t}|s_{t})$ based on Eq. (3), and optimizing the parameter $\theta$ based on Eq. (2). This line of work provides many insights that have been used to derive modern IRL (Fu et al., 2018; Reddy et al., 2019; Gao et al., 2018; Garg et al., 2021) and RL (Haarnoja et al., 2017, 2018) algorithms. Unlike these prior works, we consider the setting where we no longer have access to the expert policy $\pi_{E}$, and instead are given a heterogeneous set of potentially suboptimal demonstrations from which we want to learn a well performing policy. To accomplish this, we first describe the model we use to capture the suboptimality of demonstrators in Section 3, following which we frame this problem within the IRL framework and derive our own method IRLEED in Section 4. Finally, we describe our experimental results in Section 5, and conclude our work in Section 6.
3 Suboptimal Demonstrator Model
In this work, we consider the case where dataset $\mathcal{D}=\{(i,\mathcal{D}_{i})\}_{i=1}^{N}$ consists of a mixture of $N$ demonstrations with varying quality. Each demonstration $\mathcal{D}_{i}=\{\tau_{j}\}_{j=1}^{M_{i}}$ contains a set of $M_{i}$ trajectories sampled from a distinct, fixed policy $\pi_{i}$. Rather than considering $\mathcal{D}$ as a homogeneous set derived from one expert policy $\pi_{E}$, our approach focuses on leveraging the specific source of each demonstration to enhance the IRL framework. By identifying the varying qualities within this mixed collection of demonstrations, we aim to more accurately deduce the ground truth policy $\pi_{E}$, as opposed to a simplistic method that averages behaviors. To achieve this, it is essential to develop a model for the demonstrators $\pi_{i}$ that accurately reflects their suboptimal behavior relative to the optimal policy $\pi_{E}$.
A popular framework for modeling decision making is the Boltzmann rationality model, which is widely used in psychology (Baker et al., 2009), economics (Luce, 1959) and RL (Laidlaw & Dragan, 2021). In this model, the likelihood of an agent choosing a particular action is proportional to the exponential of the reward associated with that action: $\pi(a|s;r,\beta)\propto\exp(\beta r(s,a))$, where the inverse temperature parameter $\beta$ controls the randomness of the choice. While this formulation provides an intuitive way to capture one form of suboptimality through $\beta$, it traditionally assumes that agents have unbiased knowledge about the true reward $r(s,a)$. To broaden this model for suboptimal demonstrations, we propose considering agents that follow a biased reward $r(s,a)+\epsilon(s,a)$, introducing $\epsilon$ as the deviation from the true reward. This modification enables us to represent an agent’s accuracy through the deviation $\epsilon$, adding a bias to their perceived rewards, and their precision through parameter $\beta$, adding variance to their action choices.
Within the context of the IRL framework described in Section 2, we recall that the expert policy $\pi_{E}$ aligns with the true reward, formulated as $r(s,a)=\theta^{\star\top}f(s,a)$. Applying the aforementioned concept, we can model a demonstrator $i$’s perceived reward as a deviation from the true reward: $r_{i}(s,a)=(\theta^{\star}+\epsilon_{i})^{\top}f(s,a)$, where $\epsilon_{i}\in\mathbb{R}^{k}$. By incorporating this altered reward $r_{i}$ into the soft Bellman policy defined in Eq. (3), we derive a recursive parameterization for demonstrator $i$’s policy:
[…]
Our model captures two distinct aspects of suboptimal behavior: (1) $\epsilon_{i}$ quantifies the demonstrator’s accuracy in estimating the true reward $r$, where $\epsilon_{i}^{\top}f(\tau)$ represents the estimation error. (2) $\beta_{i}$ quantifies the demonstrator’s precision in action selection, where $\beta_{i}\rightarrow\infty$ interpolates from the soft Bellman policy, which samples actions according to their Q–values, to the standard Bellman policy, which chooses actions that maximize the Q–value. This model offers a versatile and mathematically convenient way to depict suboptimal demonstrator behaviors within the soft Bellman policy domain, as well as extend the results of Maximum Causal Entropy IRL to the suboptimal setting.
Remark 3.1.
Before moving forward, it is important to note that this model extends outside the scope of problems where the true reward $r$ is an affine transformation of some feature vector $f$. In the general setting, we can parameterize demonstrator $i$’s reward as $r_{i}=r+\epsilon_{i}$, a combination of the true reward $r\in\mathcal{R}$ and their deviation $\epsilon_{i}\in\mathcal{R}$, where $r$ and $\epsilon_{i}$ can be represented by neural networks to allow for general function approximation. We opt to formulate our proposed model within the classic feature matching scope, as this facilitates direct comparison with the seminal works used to derive our results in the following section. However, our proposed model’s applicability extends beyond this traditional framework, as we elaborate in Section 4.3, and demonstrate through experiments in Section 5. Until then, we utilize the feature matching framework described, referencing it as standard IRL.
4 Learning from Suboptimal Data with IRL
Up to this point, we have formulated a model for each demonstrator $i$ based on the parameter $\theta^{\star}$, which defines the true reward, and parameters $\epsilon_{i}$ and $\beta_{i}$, which characterize demonstrator suboptimalities. However, we do not have access to any of these parameters, and are instead provided with a dataset of demonstrations $\mathcal{D}$. Recall that standard IRL assumes that all demonstrations come from the optimal policy $\pi_{E}$, and follows the iterative approach described in Section 2 to find it. However, applying this approach to the heterogeneous regime will yield an averaged policy, which can lead to subpar performance if a significant portion of the provided demonstrations are suboptimal.
To refine this, we introduce IRLEED, an innovative extension of the IRL framework tailored for the suboptimal setting. Firstly, Section 4.1 details how IRLEED extends the standard IRL framework, which corresponds to the predefined feature matching setting. Following this, Section 4.2 presents insightful comparisons between our IRLEED approach and the standard IRL framework, along with the related behavior cloning approach, ILEED. Finally, Section 4.3 explains how IRLEED can be applied to the generalized IRL setting, addressing the practical considerations necessary for implementing our framework in conjunction with existing IRL methodologies.
4.1 IRLEED
Utilizing the dataset $\mathcal{D}=\{(i,\mathcal{D}_{i})\}_{i=1}^{N}$, we can jointly estimate the unknown parameters by maximizing the likelihood defined below:
$\mathcal{L}(\theta,\epsilon,\beta)=\log\prod_{\mathcal{D}_{i}\in\mathcal{D}}\prod_{\tau_{j}\in\mathcal{D}_{i}}P(\tau_{j}|\theta,\epsilon_{i},\beta_{i}),$ (7)
where $\epsilon=\{\epsilon_{i}\}_{i=1}^{N}$, $\beta=\{\beta_{i}\}_{i=1}^{N}$, and the probability $P(\tau_{j}|\theta,\epsilon_{i},\beta_{i})$ is given by:
$P(\tau|\theta,\epsilon_{i},\beta_{i})=p_{0}(s_{0})\prod_{t=0}^{T}\pi_{\theta,\epsilon_{i},\beta_{i}}(a_{t}|s_{t})\mathcal{T}(s_{t+1}|s_{t},a_{t}).$ (8)
Under the specified demonstrator model, this likelihood serves as a suitable loss function. Specifically, rewriting the joint likelihood equivalently only over $\theta$:
$\mathcal{L}(\theta)=\max_{\epsilon,\beta}\mathcal{L}(\theta,\epsilon,\beta),$ (9)
it is easy to show that $\mathcal{L}(\theta)$ is a proper objective function.
Proposition 4.1.
$\theta^{\star}$ is the (non–unique) maximizer of $\mathcal{L}(\theta)$.
Moreover, this formulation provides us with a convenient way to derive the gradient of likelihood function $\mathcal{L}(\theta,\epsilon,\beta)$.
Lemma 4.2.
Given the likelihood $\mathcal{L}$ defined according to Eq. (7) and demonstrator policies $\pi_{\theta,\epsilon_{i},\beta_{i}}$ parameterized by Eq. (6), we can compute the gradients of $\mathcal{L}$ as follows:
$\nabla_{\theta}\mathcal{L}$ $=\sum_{i=1}^{N}\beta_{i}(\tilde{f}_{\mathcal{D}_{i}}-\bar{f}_{\pi_{\theta,\epsilon_{i},\beta_{i}}}),$ (10)
$\nabla_{\epsilon_{i}}\mathcal{L}$ $=\beta_{i}(\tilde{f}_{\mathcal{D}_{i}}-\bar{f}_{\pi_{\theta,\epsilon_{i},\beta_{i}}}),$ (11)
$\frac{\partial\mathcal{L}}{\partial\beta_{i}}$ $=(\theta+\epsilon_{i})^{\top}(\tilde{f}_{\mathcal{D}_{i}}-\bar{f}_{\pi_{\theta,\epsilon_{i},\beta_{i}}}).$ (12)
Proof.
First, note that parameter $\beta_{i}$ can be folded inside the modified soft Bellman policy of Eq. (6) as the reward estimate: $\beta_{i}(\theta+\epsilon_{i})^{\top}f(s_{t},a_{t})$. Second, we can express the likelihood as $\mathcal{L}=\sum_{i=1}^{N}\mathcal{L}_{i}$, where $\mathcal{L}_{i}=\log\prod_{\tau_{j}\in\mathcal{D}_{i}}P(\tau_{j}|\theta,\epsilon_{i},\beta_{i})$. Setting $u_{i}=\beta_{i}(\theta+\epsilon_{i})$, we know that $\nabla_{u_{i}}\mathcal{L}_{i}=\tilde{f}_{\mathcal{D}_{i}}-\bar{f}_{\pi_{\theta,\epsilon_{i},\beta_{i}}}$ $\forall i\in\{1,\ldots,N\}$ (see (Ziebart, 2010), Theorem 6.2 and Lemma A.2). Applying the chain rule completes the proof. ∎
Notably, this result has an intuitive interpretation. Recall that for each demonstrator $i$, $\tilde{f}_{\mathcal{D}_{i}}-\bar{f}_{\pi_{\theta,\epsilon_{i},\beta_{i}}}$ represents the difference between the empirical feature vector given samples from their policy, $\tilde{f}_{\mathcal{D}_{i}}$, and the expected feature vector under the probabilistic model, $\bar{f}_{\pi_{\theta,\epsilon_{i},\beta_{i}}}$. With this in mind, Eq. (11) shows that we update $\epsilon_{i}$ to directly match feature expectation with demonstrator $\pi_{i}$, just like in standard IRL. On the other hand, Eq. (10) shows that we update $\theta$ to match a weighted average over feature expectations provided by all demonstrators, where their respective precision $\beta_{i}$ determines their contribution. Finally, Eq. (12) shows that we update this precision $\beta_{i}$ to balance the expected returns of the demonstrator, $\pi_{i}$, and our probabilistic model, $\pi_{\theta,\epsilon_{i},\beta_{i}}$, under our estimate of the perceived reward, $r_{i}(s,a)=(\theta+\epsilon_{i})^{\top}f(s,a)$. This way, when the probabilistic model outperforms the demonstration set under the estimated reward $r_{i}$, we decrease the precision $\beta_{i}$ to lower its relative performance, and vice versa.
The above formulation gives us a general algorithm that extends the IRL framework described in Section 2. Specifically, IRLEED follows an iterative approach to finding the true reward parameter $\theta^{\star}$: given demonstration set $\mathcal{D}=\{(i,\mathcal{D}_{i})\}_{i=1}^{N}$ and estimates $\theta,\epsilon,\beta$, for each demonstrator $i\in\{1,\ldots,N\}$, alternate between inferring the policy $\pi_{\theta,\epsilon_{i},\beta_{i}}(a_{t}|s_{t})$ based on Eq. (6), and optimizing the parameters $\theta,\epsilon_{i},\beta_{i}$ based on Eqs. (10), (11), (12), respectively.
As aforementioned, many different techniques can be used for both the inference and optimization procedures. For example, to infer the policy $\pi_{\theta,\epsilon_{i},\beta_{i}}(a_{t}|s_{t})$ given $\theta,\epsilon_{i},\beta_{i}$, one can use soft versions of the standard value iteration and Q-learning algorithms depending on whether the dynamics $\mathcal{T}$ are provided or not (Bloem & Bambos, 2014). To compute feature expectations $\bar{f}_{\pi_{\theta,\epsilon_{i},\beta_{i}}}$ based on the inferred policy, one can directly compute the expectations by utilizing the standard dynamic programming operator when the dynamics $\mathcal{T}$ are known, or estimate the expectations by utilizing Monte Carlo simulations of the MDP when the dynamics $\mathcal{T}$ are unknown.
4.2 Comparisons to Standard IRL and ILEED
We can see that this likelihood maximization problem is closely related to the maximum causal entropy IRL framework described in Eqs. (1)- (5). Specifically, IRLEED makes two modification: (1) For each $i\in\{1,\ldots,N\}$, it defines individual dual variables $\theta_{i}=\theta+\epsilon_{i}$ that correspond to feature matching constraint $\tilde{f}_{\mathcal{D}_{i}}=\bar{f}_{\pi_{\theta,\epsilon_{i},\beta_{i}}}$, tying the demonstrator policies together. (2) It uses a learnable parameter $\beta_{i}$ to tune the magnitude of uncertainty related to demonstrations provided in $\mathcal{D}_{i}$. Moreover, IRLEED generalizes the standard IRL framework.
Remark 4.3.
IRLEED recovers the standard IRL framework when we set $\epsilon_{i}=[0]^{k}$ and $\beta_{i}=1$ as constants.
To show the advantage of IRLEED within the suboptimal demonstration setting, we note that the policy recovered by standard IRL can only perform as well as the average demonstration $\mathcal{D}_{i}$ provided. This is a direct result of the feature matching constraint imposed by the IRL objective.
Proposition 4.4.
Given demonstrations $\mathcal{D}=\{(i,\mathcal{D}_{i})\}_{i=1}^{N}$ produced by suboptimal policies $\pi_{i}$ defined according to Eq. (6), let $\theta_{IRL}$ denote the naive solution to the standard IRL problem defined by Eq. 2, which treats demonstration set $\mathcal{D}$ as one homogeneous set produced by $\pi_{E}$. The performance of IRL is bounded by:
$\operatorname{\mathbb{E}}_{\pi_{\theta_{IRL}}}[r(s,a)]=\frac{1}{N}\sum_{i=1}^{N}\operatorname{\mathbb{E}}_{\pi_{i}}[r(s,a)]\leq\operatorname{\mathbb{E}}_{\pi_{\theta^{\star}}}[r(s,a)],$ (13)
where for simplicity, we assume that all demonstrators provide the same number of trajectories $M_{i}=M$.
As mentioned in Proposition 4.1, IRLEED can recover the true reward parameter $\theta^{\star}$, which outperforms $\theta_{IRL}$ unless the demonstrations provided are optimal. Unfortunately, since $\theta^{\star}$ is not the unique solution to IRLEED, we can not make guarantees on recovering the true reward parameter $\theta^{\star}$. Nonetheless, our experimental results in Section 5 demonstrate that IRLEED improves performance over IRL frameworks in both the feature matching setting described here, as well as more complex scenarios which include high dimensional input spaces and human demonstrations. Finally, we note that when the MDP is deterministic, we can express the policy without recursion:
$\pi_{\theta,\epsilon_{i},\beta_{i}}(\tau)\propto\exp(\beta_{i}(\theta+\epsilon_{i})^{\top}f(\tau)),$ (14)
where $f(\tau)=\sum_{(s,a)\in\tau}f(s,a)$, and $\pi(\tau)$ can directly map from trajectories due to the deterministic dynamics $\mathcal{T}$. Replacing the probability $P(\tau|\theta,\epsilon_{i},\beta_{i})$ with $\pi_{\theta,\epsilon_{i},\beta_{i}}(\tau)$, we can see that maximizing the likelihood defined by Eq. (7) corresponds to the supervised learning approach employed by ILEED (Beliaev et al., 2022).
Remark 4.5.
IRLEED recovers the ILEED framework if we assume all demonstrators $i$ have full knowledge of the true reward, $\epsilon_{i}=[0]^{k}$, and ignore the underlying dynamics of the MDP.
Hence unlike ILEED, our framework provides us with a dynamics–aware solution to the problem of learning from mixtures of suboptimal demonstrations. As we demonstrate in Section 5, IRLEED can be utilized offline to achieve superior performance compared to ILEED.
4.3 Extension to Generalized IRL
While the results in the previous sections provide insight on how to learn from suboptimal demonstrations using IRL, relying on hand-crafted features $f$ can hinder our ability to express complex behavior. In general, the true reward $r$ can belong to a non restrictive set of functions =$\mathcal{R}=\{r:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}\}$. In this case, one can utilize the generalized IRL objective instead:
$\max_{r\in\mathcal{R}}\min_{\pi\in\Pi}\operatorname{\mathbb{E}}_{\pi_{E}}[r(s,a)]-\operatorname{\mathbb{E}}_{\pi}[r(s,a)]-H(\pi)-\psi(r),$ (15)
where $\psi$ is a convex reward regularizer that smoothly penalizes differences between occupancy measures to prevent overfitting (Ho & Ermon, 2016). A naive solution to this nested min-max objective involves an outer loop learning the reward, and an inner loop executing maximum entropy RL with this reward to find the optimal policy. As before, the optimal policy under reward $r$ takes the form of the soft Bellman policy defined in Equation 3, where $r(s,a)$ replaces the linear reward $\theta^{\top}f(s,a)$. The description of IRLEED in the generalized setting follows directly.
Since the suboptimal demonstrator model defined in Equation 6 extends to this setting, we can parameterize our estimate of demonstrator $i$’s policy with the soft Bellman policy $\hat{\pi}_{i}$ under the reward estimate $r_{i}=r+\epsilon_{i}$ and precision $\beta_{i}$. Utilizing the dataset $\mathcal{D}=\{(i,\mathcal{D}_{i})\}_{i=1}^{N}$, the outer loop of the generalized IRLEED objective is simplified to:
$\max_{r,\epsilon,\beta}\sum_{i=1}^{N}\beta_{i}\big{(}\operatorname{\mathbb{E}}_{\mathcal{D}_{i}}[r_{i}(s,a)]-\operatorname{\mathbb{E}}_{\hat{\pi}_{i}}[r_{i}(s,a)]\big{)}-\psi(\beta_{i}r_{i}),$ (16)
where in practice, the unknown functions $r$ and $\epsilon=\{\epsilon_{i}\}_{i=1}^{N}$ can be parameterized by neural networks. Note that under a constant reward regularizer $\psi$, if we express the rewards as $r_{i}(s,a)=(\theta+\epsilon_{i})^{\top}f(s,a)$, then the gradients of the above objective are identical to the ones described in Lemma 4.2.
In summary, we have shown how any learning algorithm that solves the generalized IRL objective defined in Equation 15 can be utilized alongside our framework by re-paremetrizing the reward and policy accordingly. Our experiment results in Section 5 demonstrate that IRLEED performs well in this generalized setting, and can even be trained offline when paired with Inverse soft-Q learning (IQ), an approach that avoids the iterative process defined by learning a single Q-function, implicitly representing both reward and policy (Garg et al., 2021).
5 Experiments
In this section, we empirically evaluate how IRLEED performs when learning from suboptimal demonstrations, using experiments in both online and offline IL settings, with simulated and human-generated data. In total, we performed three sets of experiments, comparing IRLEED to maximum entropy IRL (Ziebart et al., 2010), Inverse soft-Q learning (IQ) (Garg et al., 2021), and ILEED (Beliaev et al., 2022). We go over the results of each experiment below, briefly detailing the corresponding setup for each. For further implementation details refer to Appendix A.
5.1 Gridworld
For our first experiment, we used a custom gridworld environment, simulating suboptimal demonstrations with varying levels of precision and accuracy. Using this environment, we compared the performance of IRLEED against maximum entropy IRL.
Setup The reward in this setting was linearly parameterized as $r(s)=\theta^{\top}f(s)$, where $f(s)$ is simply a one hot encoding for each gridworld state. To implement IRL, we utilized stochastic value iteration on the reward estimate $\theta$ to infer the policy $\pi$, and Monte Carlo simulations of the MDP to estimate the feature expectations required to optimize $\theta$, repeating both steps until convergence. We used the same processes to implement IRLEED according to the procedure outlined in Section 4, updating $\epsilon$ and $\beta$ only after $\theta$ converged, and repeating this for two iterations.
For each level of precision and accuracy, we randomly generated $5$ suboptimal policies according to Equation 6, where $\beta_{i}$ was sampled from a uniform distribution starting at 00, with mean equal to the precision level, and $\epsilon_{i}$ was sampled from a multivariate normal distribution $\mathcal{N}([0]^{k},\mathbf{I}_{k}/\lambda^{2})$, with $\lambda$ equal to the accuracy level. We tested a total of $121$ dataset settings, collecting $40$ trajectories from each policy to compose our dataset, and comparing the performance of IRLEED and IRL over $100$ seeds for each setting.
Results We first compare how IRLEED and IRL recover the reward $r(s,a)$ when provided with suboptimal demonstrations. We visualize the averaged result for one dataset setting in Figure 2, which shows the true reward, the dataset quality, and the recovered rewards under both methods. We can see that the provided demonstrations are misaligned with the ground truth reward: the state visitation frequency for the top left corner is higher due to demonstrator suboptimalities. As expected, the feature matching constraint of IRL absorbs this suboptimality. Although the reward recovered by IRL contains information about the true reward, it is now incorrectly biasing the top left corner. On the other hand, IRLEED is able to remove this bias, providing us with a better estimate of the ground truth reward.
To see the effect of dataset quality, we use Figure 3 to show the relative improvement of the policy recovered by IRLEED over maximum entropy IRL for each dataset setting. We can see from the left plot that IRLEED provides more improvement over IRL as we decrease accuracy $\epsilon$ (by decreasing $\lambda$), or increase precision $\beta$. This signifies the importance of modeling the precision $\epsilon$, or reward bias, in addition to the accuracy $\beta$, or action variance, when learning from suboptimal demonstrations. This point is further highlighted in the right plot, which shows that although both IRLEED and IRL suffer when demonstrator accuracy is diminished, IRLEED can tolerate lower demonstrator accuracy compared to IRL. In total, IRLEED provided a $30.3\%$ improvement over IRL for the $121$ settings tested.
Figure 2: We visualize the reward recovered by IRLEED and IRL when trained using suboptimal demonstrations. Top left shows the true reward, where the three yellow corners are terminal states. Top right shows the normalized state visitation frequency over the entire dataset. Bottom left and right show the normalized rewards recovered by IRLEED and IRL respectively.
Figure 3: We compare the performance of the policies recovered by IRLEED and IRL. The left plot shows the relative performance of IRLEED over IRL under varying dataset settings, where the top right corner corresponds to expert data. The right plot shows the performance of both policies as we increase the average accuracy $\epsilon$ of demonstrators in the dataset, corresponding to the data outlined in red in the left plot.
5.2 Simulated Control Tasks
For our second experiment, we used two control tasks, Cartpole, and Lunar Lander, simulating suboptimal demonstrations with varying levels of expertise and accuracy. Using these two environments, we compared the performance of IRLEED, IQ, and ILEED in both the online and offline settings.
Setup To implement IQ, and setup the training and evaluation procedures, we utilized the codebase provided by the authors. We implemented ILEED alongside this codebase, using the training and evaluation procedures. For IRLEED, we built on top of the aforementioned IQ implementation by adding a learnable parameter $\beta_{i}$, and an additional critic network $\epsilon_{i}$, for each demonstrator $i$. Since IQ learns a reward and policy directly by using a single Q function as a critic, we used our additional critic network $\epsilon_{i}$ to directly add bias to the state action function instead of the reward.
For this experiment, we used $4$ dataset settings corresponding to the precision and accuracy levels: (high, high), (high, low), (low, high), and (low, low). For each setting, we randomly generated $3$ suboptimal policies according to $\pi_{i}(a|s)\propto\exp(\beta_{i}(Q(s,a)+\alpha\epsilon_{i}(s,a)))$, where $Q(s,a)$ is a pretrained critic network, $\epsilon_{i}(s,a)$ is a randomly initialized critic network, and $\alpha$ is a scaling factor which is set to zero for the high accuracy setting. For the low precision setting, $\beta_{i}$ was sampled from a uniform distribution, whereas for the high precision setting, the stochasticity of the policy was removed by following the maximum state action value. We ran this experiment in both the online and offline setting, comparing the performance of IRLEED, IQ, and ILEED, using $30$ seeds for each dataset setting.
Results We first compare how IRLEED improves over IQ, with online access to environment interactions. We provide these results in Tables 1 and 2, which show the return of the recovered policy relative to the return of the best demonstrator in that dataset. We can see that IRLEED outperforms its ablated counterpart, IQ, under all $3$ suboptimal dataset settings (not including the clean dataset), for both environments. Furthermore, we see that in the more complex Lunar Lander environment, IQ can not perform as well as the best demonstrator, whereas IRLEED consistently performs on par or better than the best demonstrator.
To have a fair comparison between IRLEED and its behaviour cloning counterpart, ILEED, we repeat these experiments without online access to environment interactions. Firstly, we can see that from Tables 1 and 2 that offline IRLEED outperforms ILEED, under all $3$ suboptimal dataset settings, for both environments. We can see that although ILEED has comparable performance in the simpler Cartpole environment, it performs poorly in the more complex Lunar Lander environment. Moreover, we can see from Table 2 that ILEED shows diminished performance when demonstrators have low accuracy $\epsilon$, signifying that it can not account for the reward biases present in these settings.
Table 1: Cartpole
We list the average return of the recovered policies (IQ, IRLEED, ILEED), relative to the return of the best demonstrator. This is provided for $4$ dataset settings (Prec., Acc.), where we list the mean performance of all demonstrators relative to the best demonstrator (Mean). The last row (H,H) corresponds to a clean dataset. The return of the best demonstrator corresponding to each setting (top to bottom) is: $260,500,411,500$. Setting Online Offline Prec. Acc. Mean IQ IRLEED ILEED IQ IRLEED L L $0.58$ $0.97$ 1.55 $1.38$ $0.96$ 1.56 L H $0.70$ $0.92$ 1.00 $1.00$ $0.96$ 1.00 H L $0.64$ $1.03$ 1.09 $1.05$ $0.95$ 1.05 H H $1.00$ $0.99$ $1.00$ $0.92$ $1.00$ $1.00$
Table 2: Lunar Lander
See Table 1. The return of the best demonstrator corresponding to each setting (top to bottom) is: $171,213,166,277$. Setting Online Offline Prec. Acc. Mean IQ IRLEED ILEED IQ IRLEED L L $0.46$ $0.68$ 1.28 $0.84$ $0.81$ 1.34 L H $0.54$ $0.84$ 1.14 $1.06$ $1.05$ 1.18 H L $0.57$ $0.94$ 1.30 $0.85$ $0.99$ 1.14 H H $0.86$ $0.93$ 0.98 $1.01$ $0.99$ $0.98$
5.3 Atari with Human Data
For our final experiment, we used two Atari environments, Space Invaders, and Qbert, utilizing both simulated and human demonstrations. Using these two environments, we compared the performance of IRLEED, IQ, and ILEED in the offline setting.
Setup Our demonstrations were comprised of two unique sources, dataset A: simulated data using a pretrained expert policy and dataset B: collected data using adept human players (Kurin et al., 2017). Since the human dataset B was acquired using screenshots of web-browsers, there is a small mismatch in sampling frequencies and colors when compared to the simulated environment used in dataset A. This provides a unique challenge akin to crowd-sourcing data from varying sources, where the goal is to see if we can achieve improved performance when training on the combined dataset AB as opposed to dataset A alone. Note that we followed the same steps outlined in the previous Section 5.2 to implement the three methods and setup the training and evaluation procedures. The results shown are averaged over $5$ seeds.
Results With this setup, we would like to test if IRLEED can improve the crowd-sourcing capability of IQ, and compare its performance to ILEED. We provide these results for both environments in Table 3, which lists the return of the recovered policies. Notably, all methods do poorly when trained exclusively on dataset B, due to the mismatch between the simulation environment and the data. On the other hand, utilizing dataset A is enough to reach decent performance. However, we can see that IRLEED is the only method that significantly improves its performance when utilizing the combined dataset AB as opposed to dataset A.
Table 3: Utilizing Demonstrations from Varying Sources
We list the average return of the recovered policies (IQ, IRLEED, ILEED). This is provided for $3$ dataset settings, simulated dataset A, human dataset B (Kurin et al., 2017), and their combination AB. Overall we see IRLEED shows a $16\%$ improvement over IQ, and a $41\%$ improvement over ILEED in the combined AB setting.
Dataset Space Invaders Qbert
ILEED IQ IRLEED ILEED IQ IRLEED
AB $802$ $755$ 910 $5511$ $8248$ 9336
A $783$ $754$ $768$ $5926$ $8092$ $8049$
B $110$ $179$ $123$ 00 $809$ $1087$
6 Discussion
In summary, this paper addresses the challenges of leveraging suboptimal and heterogeneous demonstrations in IL by introducing IRLEED, a novel IRL framework. By integrating a general model of demonstrator suboptimality that accounts for reward bias and action variance, with a maximum entropy IRL framework, IRLEED successfully mitigates the limitations inherent in standard IL algorithms, which often fail to account for the diversity and imperfections in real-world data. We empirically demonstrate that IRLEED can be used alongside existing IRL techniques to improve their performance in the presence of suboptimal and heterogeneous demonstrations. Moreover, we show that IRLEED outperforms its behavior cloning counterpart ILEED in the offline setting, demonstrating the importance of utilizing a dynamics-aware approach that considers reward bias. Although we have shown that IRLEED improves over IRL in a multitude of settings, we have not provided guarantees that state whether IRLEED will recover a better reward estimate compared to IRL. One direction for future work could be to provide more theoretical analysis comparing IRLEED to IRL under simpler settings with strict assumptions.
7 Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Implementation Details
Below we detail the experimental setups utilized in Section 5, and provide information on the computational resources used for our work. Note that we provide a working implementation of the Gridworld experiments found in Section 5.1 as part of our supplementary material, and plan to release a public implementation of IRLEED with all experiments and datasets used after publication.
A.1 Gridworld
To implement both IRLEED and IRL for the Gridworld experiments in Section 5.1, we utilized stochastic value iteration on the reward estimate $\theta$ to infer the policy $\pi$, and Monte Carlo simulations of the MDP to estimate the feature expectations required to optimize $\theta$, repeating both steps until convergence. As stated, when using IRLEED, we only updated $\epsilon$ and $\beta$ after $\theta$ converged, and repeated this for two iterations. For stochastic value iteration, we used a discount factor of $0.9$, a maximum horizon of $100$ timesteps, and a convergence criteria of $41e-$ 4 on the inferred state action values. For Monte Carlo simulations, we used $100$ episode samples. For optimization, we used stochastic gradient ascent with learning rates of $0.05,0.1,0.2$ for $\beta$, $\epsilon$, $\theta$, respectively, and a convergence criteria of $41e-$ 4 on $\theta$. For IRL, we initialized $\theta=[0.1]^{k}$, and set $\beta_{i}=1$ and $\epsilon_{i}=[0]^{k}$ to constants for all demonstrators, removing their effect. For evaluation, we used the policy recovered from the final estimate of $\theta$, sampling $100$ episodes to measure the mean reward.
To create datasets for varying levels of precision and accuracy, we randomly generated $5$ suboptimal policies according to Equation 6, where $\beta_{i}$ was sampled from a uniform distribution starting at 00, with mean equal to the precision level, and $\epsilon_{i}$ was sampled from a multivariate normal distribution $\mathcal{N}([0]^{k},\mathbf{I}_{k}/\lambda^{2})$, with $\lambda$ equal to the accuracy level. For $\beta$, the maximum values used for the uniform distribution were: $0.4,0.5,1,1.5,2,2.5,3,3.5,4,4.5,5$. For $\lambda$, the values used were: $2,2.5,3,3.5,4,4.5,5,5.5,6,10,\infty$, where $\infty$ corresponds to setting $\epsilon_{i}=[0]^{k}$ to remove its effect. We tested a total of $121$ dataset settings, collecting $40$ trajectories from each policy to compose our dataset, and comparing the performance of IRLEED and IRL over $100$ seeds for each setting.
A.2 Simulated Control Tasks
To implement IQ and setup the training and evaluation procedures for the simulated control tasks in Section 5.2, we utilized the codebase provided by the authors. [^2] https://github.com/Div99/IQ-Learn. For all experiments, we utilized the provided hyperparameters for IQ. To implement ILEED, we followed the codebase provided by the authors, implementing the modified behaviour cloning loss alongside the IQ implementation. [^3] https://github.com/Stanford-ILIAD/ILEED For IRLEED, we built on top of the aforementioned IQ implementation by adding a learnable parameter $\beta_{i}$, and an additional critic network $\epsilon_{i}$, for each demonstrator $i$. This way, we sampled suboptimal policies according to: $\pi_{i}(a|s)\propto\exp(\beta_{i}(Q(s,a)+\alpha\epsilon_{i}(s,a))$. All three methods utilized the same architecture for the critic networks, and followed the original hyperparamters used for the IQ implementation. For ILEED, the critic network was used to compute action probabilities using a final softmax layer. For IRLEED and ILEED, $\beta_{i}$ was initialized to the same inverse temperature parameter used in the IQ implementation, and the learning rate was set to $45e-$ 4. For IRLEED, the additional critic networks $\epsilon_{i}(s,a)$ were initialized with the same procedure as the critic network $Q(s,a)$, utilized the same learning rate divided by $10$, and utilized the scaling factor $\alpha=0.01$. We trained all methods using $50,000$ and $100,000$ time steps for Cartpole and Lunar Lander, respectively. For evaluation, we utilized the final policy, sampling $300$ episodes to measure the mean reward.
To create datasets for varying levels of precision and accuracy, we randomly generated $3$ suboptimal policies according to $\pi_{i}(a|s)\propto\exp(\beta_{i}(Q(s,a)+\alpha\epsilon_{i}(s,a))$, where $Q(s,a)$ was a pretrained critic network provided by the original IQ implementation, and $\epsilon_{i}(s,a)$ was a randomly initialized critic network. For Cartpole, we utilized $\alpha=0$ and $\alpha=20$ for the high and low accuracy settings, respectively. For Lunar Lander, we utilized $\alpha=0$ and $\alpha=0.2$ for the high and low accuracy settings, respectively. For the low precision setting, $\beta_{i}$ was sampled from a uniform distribution between 00 and $1$ for Cartpole, and between 00 and $100$ for Lunar Lander. For the high precision setting, the effect of $\beta$ was removed and the policies were sampled according to the maximum state action values: $\pi_{i}(a|s)=\max(Q(s,a)+\alpha\epsilon_{i}(s,a))$. We ran this experiment in both the online and offline setting, comparing the performance of IRLEED, IQ, and ILEED, using $30$ seeds for each dataset setting.
A.3 Atari with Human Data
As aforementioned, the implementation for the Atari experiments in Section 5.3 was identical to the one described above in Section A.2. For training, we utilized $1,000,000$ time steps for all methods. For IRLEED, we did not update the error critic networks $\epsilon_{i}(s,a)$ or the temperature parameters $\beta_{i}$ for the first $300,000$ timesteps. For evaluation, we sampled $10$ episodes from the policies recovered during the last $5$ epochs of training, where each epoch corresponded to $5,000$ time steps.
The two datasets used are publicly available, where dataset A: simulated data using a pretrained expert policy, was provided alongside the original IQ implementation (Garg et al., 2021) and dataset B: collected data using adept human players, was part of a larger human study (Kurin et al., 2017). For Space Invaders, dataset A contained $5$ trajectories with an average return of $1285$, whereas dataset B contained $5$ trajectories with an average return of $1801$. For Qbert, dataset A contained $10$ trajectories with an average return of $14760$, whereas dataset B contained $10$ trajectories with an average return of $17340$. We ran this experiments using the three combinations of these datasets, averaging our results over $5$ seeds.
A.4 Computational Resources
All of the experiments were performed on an internal cluster containing the 20C/40T Intel Xeon Silver 4114 CPU, 64GB RAM, and 4×GTX-1080 GPUs. Obtaining results for the Gridworld experiment with $121$ datasets, each with $100$ seeds, took roughly $3$ days. Obtaining results for the simulated control tasks, Cartpole and Lunar Lander, took roughly $4$ days each. Training models on the Atari environments took $40$ hours per seed for IQ and IRLEED, and $20$ hours per seed for ILEED.

### v2

@@
 However, standard IL algorithms consider these datasets as homogeneous, thereby inheriting the deficiencies of suboptimal demonstrators.
-Previous approaches to this issue typically rely on impractical assumptions like high-quality data subsets, confidence rankings, or explicit environmental knowledge.
+Previous approaches to this issue rely on impractical assumptions like high-quality data subsets, confidence rankings, or explicit environmental knowledge.
 This paper introduces IRLEED, Inverse Reinforcement Learning by Estimating Expertise of Demonstrators, a novel framework that overcomes these hurdles without prior knowledge of demonstrator expertise.
@@
 Experiments in both online and offline IL settings, with simulated and human-generated data, demonstrate IRLEED’s adaptability and effectiveness, making it a versatile solution for learning from suboptimal demonstrations.
-Learning from Suboptimal Demonstrations, Inverse Reinforcement Learning
-Figure 1: IRLEED can be applied in the suboptimal demonstration setting to estimate the ground truth reward $r$, which is used to find the optimal policy.
-Left: A heterogeneous dataset is collected from multiple sources with varying degrees of optimality.
+Codehttps://github.com/mbeliaev1/IRLEED
+Extended versionhttps://arxiv.org/abs/2402.01886
+Figure 1: IRLEED is applied in the suboptimal setting to estimate the ground truth reward $r$, which is used to find the optimal policy.
+Left: A heterogeneous dataset is collected from multiple sources with varying optimality.
 We categorize this optimality by using accuracy to represent the reward bias, and precision to represent the variance in action choices.
-Right: We can infer the demonstrator policies using a model for suboptimal behaviour based on the Boltzmann rationality principle, which captures both the accuracy $\epsilon_{i}$, and the precision $\beta_{i}$, as compared to the ground truth reward $r$.
-Middle: Using estimates of the demonstrator policies $\hat{\pi}_{i}$ along with the demonstrations $\mathcal{D}_{i}$, we can optimize for the true reward $r$, as well as the parameters that capture accuracy $\epsilon$ and precision $\beta$.
+Right: We infer the demonstrator policies using a model for behavior based on the Boltzmann rationality principle, which captures both the accuracy $\epsilon_{i}$, and the precision $\beta_{i}$, as compared to the ground truth reward $r$.
+Middle: Using estimates of the demonstrator policies $\hat{\pi}_{i}$ along with the demonstrations $\mathcal{D}_{i}$, we can optimize for the true reward $r$, and the parameters that capture accuracy $\epsilon$ and precision $\beta$.
 1 Introduction
-Reinforcement Learning (RL) has proven to be a versatile and powerful tool across a wide range of applications, from navigating autonomous vehicles, to playing complex games like Go and Chess.
-The efficacy of RL algorithms is largely attributed to their ability to optimize hand-crafted reward functions (Brockman et al., 2016).
-However, designing these functions is often challenging and impractical in complex environments (Hadfield-Menell et al., 2017).
-A prevalent alternative to this dilemma is Imitation Learning (IL), which bypasses the need for cumbersome reward engineering by leveraging expert demonstrations to instill desired behaviors (Argall et al., 2009).
-The two main approaches utilized in IL are: behavioral cloning (Pomerleau, 1991), which acquires a policy through a supervised learning approach using state-action pairs provided by the expert, and Inverse Reinforcement Learning (IRL) (Ng & Russell, 2000), which solves for the reward function that makes the expert’s behavior optimal, subsequently facilitating the training of an IL policy.
-Despite the effectiveness of behavior cloning in simple environments with large amounts of data, learning a policy to fit single time step decisions leads to compounding errors due to covariance shift (Ross & Bagnell, 2010).
-On the other hand, IRL learns a reward function that prioritizes entire trajectories over others, considering the sequential nature of the decision making problem at hand (Abbeel & Ng, 2004).
-Consequently, the success of IRL has spurred the development of IL techniques that either explicitly or implicitly incorporate environment dynamics (Ho & Ermon, 2016; Fu et al., 2018; Kostrikov et al., 2019; Garg et al., 2021).
+Reinforcement Learning (RL) has proven to be a powerful tool across a wide range of applications, from controlling robotic systems, to playing complex games like Go and Chess.
+The efficacy of RL algorithms is largely attributed to their ability to optimize hand-crafted reward functions (Brockman et al.
+2016).
+However, designing these functions is often challenging and impractical in complex environments (Hadfield-Menell et al.
+2017).
+A prevalent alternative to this dilemma is Imitation Learning (IL), which bypasses the need for cumbersome reward engineering by leveraging expert demonstrations to instill desired behaviors (Argall et al.
+2009).
+The two main approaches utilized in IL are: behavioral cloning (Pomerleau 1991), which acquires a policy through a supervised learning approach using state-action pairs provided by the expert, and Inverse Reinforcement Learning (IRL) (Ng and Russell 2000), which solves for the reward function that makes the expert’s behavior optimal, subsequently facilitating the training of an IL policy.
+Despite the effectiveness of behavior cloning in simple environments with large amounts of data, learning a policy to fit single time step decisions leads to compounding errors due to covariance shift (Ross and Bagnell 2010).
+On the other hand, IRL learns a reward function that prioritizes entire trajectories over others, considering the sequential nature of the decision making problem at hand (Abbeel and Ng 2004).
+Consequently, the success of IRL has spurred the development of IL techniques that either explicitly or implicitly incorporate environment dynamics (Ho and Ermon 2016; Fu, Luo, and Levine 2018; Kostrikov, Nachum, and Tompson 2019; Garg et al.
+2021).
 However, a critical assumption in these methodologies is the availability of high-quality demonstration data.
-In many practical situations, especially with crowd-sourced or varied data, the quality of demonstrations is inconsistent (Mandlekar et al., 2021; Belkhale et al., 2023).
-For example, foundation models like GPT-3 which are trained on large and uncurated datasets, exhibit limitations such as generating biased outputs that negatively stereotype certain groups (Abid et al., 2021).
-In robotics, data curation becomes vital, since utilizing smaller real world datasets with noise or biases can lead to hazardous situations.
-Furthermore, assuming uniformity in demonstration quality overlooks the unique intentions of individual demonstrators, potentially leading to suboptimal learning outcomes (Eysenbach et al., 2018).
+In many practical situations, especially with crowd-sourced or varied data, the quality of demonstrations is inconsistent (Mandlekar et al.
+2021; Belkhale, Cui, and Sadigh 2023).
+For robotics, data curation becomes vital, since utilizing smaller real world datasets with noise or biases can lead to hazardous situations.
+Furthermore, assuming uniformity in demonstration quality overlooks the unique intentions of individual demonstrators, potentially leading to suboptimal learning outcomes (Eysenbach et al.
+2018).
 Therefore, it is critical to develop IL methods that can account for the heterogeneity and suboptimality of demonstration data.
-Addressing the challenges posed by learning from suboptimal and heterogeneous demonstrations is complex.
-One line of research focuses on heterogeneous demonstrations without considering their quality (Chen et al., 2020, 2023), assuming that all provided demonstrations are optimal.
-Other works consider suboptimal demonstrations, but require unrealistic prerequisites such as explicit knowledge of environment dynamics (Brown et al., 2020; Chen et al., 2021), a set of confidence scores over the demonstrations (Zhang et al., 2021; Cao & Sadigh, 2021; Wu et al., 2019; Kuhar et al., 2023), or direct access to a subset of expert demonstrations (Xu et al., 2022; Yang et al., 2021).
-A recently proposed approach that avoids these assumptions is ILEED (Beliaev et al., 2022), which directly models the expertise of demonstrators within a supervised learning framework, thus enabling one to learn from the optimal behavior while filtering out the suboptimal behavior of each demonstrator.
+Related Work Addressing the challenges posed by learning from suboptimal and heterogeneous demonstrations is complex.
+One line of research focuses on heterogeneous demonstrations without considering their quality (Chen et al.
+2020, 2023), assuming that all demonstrations are optimal.
+Other works consider suboptimal demonstrations, but require unrealistic prerequisites such as explicit knowledge of environment dynamics (Brown, Goo, and Niekum 2020; Chen, Paleja, and Gombolay 2021), a set of confidence scores over the demonstrations (Zhang et al.
+2021; Cao and Sadigh 2021; Wu et al.
+2019; Kuhar et al.
+2023), or direct access to a subset of expert demonstrations (Shiarlis, Messias, and Whiteson 2016; Xu et al.
+2022; Yang, Levine, and Nachum 2021).
+A recent approach that avoids these assumptions is ILEED (Beliaev et al.
+2022), which directly models the expertise of demonstrators within a supervised learning framework, thus enabling one to learn from the optimal behavior while filtering out the suboptimal behavior of each demonstrator.
 While this method was proven effective, it is based on a behavioral cloning formulation, which overlooks the underlying environment dynamics, and its representation of suboptimal behavior is limited, presuming demonstrators to be noisy approximations of experts.
-To overcome these challenges, this paper introduces IRLEED, Inverse Reinforcement Learning by Estimating Expertise of Demonstrators.
+Overview To overcome these challenges, this paper introduces IRLEED, Inverse Reinforcement Learning by Estimating Expertise of Demonstrators.
 As illustrated in Figure 1, IRLEED is a novel framework for IRL that accounts for demonstrator suboptimality without prior expertise knowledge.
@@
   • We provide comparative insights against standard IRL and the behavior cloning approach, ILEED, showing how IRLEED generalizes both methods.
-  • We empirically validate the success of our method against relevant baselines in both online and offline IL scenarios, utilizing both simulated and human datasets.
+  • We empirically validate the success of our method against relevant baselines in both online and offline IL scenarios.
 2 Preliminary
-Notations We consider the standard Markov decision process (MDP) setting, defined by a tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{T},p_{0},r,\gamma)$, where $\mathcal{S}$, $\mathcal{A}$ represent state and action spaces, $\mathcal{T}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ represents the dynamics, $p_{0}:\mathcal{S}\rightarrow[0,1]$ represents the initial state distribution, $r\in\mathcal{R}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ represents the reward function, and $\gamma\in(0,1)$ represents the discount factor.
-In Section 3 and Section 4, we describe our method using finite state and action spaces, $\mathcal{S}$ and $\mathcal{A}$, but our experiments later utilize continuous environments.
+Notations We use the Markov decision process (MDP) setting, defined by a tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{T},p_{0},r,\gamma)$, where $\mathcal{S}$, $\mathcal{A}$ represent state and action spaces, $\mathcal{T}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ represents the dynamics, $p_{0}:\mathcal{S}\rightarrow[0,1]$ represents the initial state distribution, $r\in\mathcal{R}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ represents the reward function, and $\gamma\in(0,1)$ represents the discount factor.
+In Sections  3 and 4, we describe our method using finite state and action spaces, $\mathcal{S}$ and $\mathcal{A}$, but our experiments later utilize continuous environments.
 We use $\pi:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ to denote a policy that assigns probabilities to actions in $\mathcal{A}$ given states in $\mathcal{S}$.
@@
 Additionally, we will use $\tilde{f}_{\mathcal{D}}=\operatorname{\mathbb{E}}_{\mathcal{D}}[f(s,a)]$ to denote the empirical expectation with respect to the trajectories $\tau_{j}=(s_{0},a_{0},\ldots,s_{T},a_{T})$ in demonstration set $\mathcal{D}=\{\tau_{j}\}_{j=1}^{M}$, where we leave out denoting trajectories of varying lengths for simplicity.
-Maximum Entropy IRL The goal of Inverse Reinforcement Learning (IRL) is to recover a reward $r\in\mathcal{R}$ that rationalizes the demonstrated behavior in dataset $\mathcal{D}$, where the provided dataset consists of trajectories $\tau$ sampled from an expert policy $\pi_{E}$.
-This general idea was first formulated as a feature-expectation matching problem (Abbeel & Ng, 2004): assuming that the feature vector $f:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}^{k}$ components quantify the expert’s behavior, to find a policy $\pi$ that performs equal to or better than the expert $\pi_{E}$, it suffices that their feature expectations match $\bar{f}_{\pi}=\bar{f}_{\pi_{E}}$.
+Maximum Entropy IRL The goal of IRL is to recover a reward $r\in\mathcal{R}$ that rationalizes the demonstrated behavior in dataset $\mathcal{D}$, where the provided dataset consists of trajectories $\tau$ sampled from an expert policy $\pi_{E}$.
+This idea was first formulated as a feature-expectation matching problem (Abbeel and Ng 2004): assuming that the feature vector $f:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}^{k}$ components quantify the expert’s behavior, to find a policy $\pi$ that performs equal to or better than the expert $\pi_{E}$, it suffices that their feature expectations match $\bar{f}_{\pi}=\bar{f}_{\pi_{E}}$.
 Unfortunately, this leads to an ill-posed problem as many policies can lead to the same feature counts.
-Maximum Entropy IRL resolves this ambiguity by choosing the policy which does not exhibit any additional preferences beyond matching feature expectations (Ziebart et al., 2008):
+Maximum Entropy IRL resolves this ambiguity by choosing the policy which does not exhibit any additional preferences beyond matching feature expectations (Ziebart et al.
+2008):
 $\operatorname*{arg\,max}_{\pi}H(\pi),\textit{ such that: }\bar{f}_{\pi}=\tilde{f}_{\mathcal{D}},$ (1)
-where in the infinite-time horizon setting, this equates to maximizing the discounted causal entropy $H(\pi)$ under the feature matching constraint (Bloem & Bambos, 2014).
+where in the infinite-time horizon setting, this equates to maximizing the discounted causal entropy $H(\pi)$ under the feature matching constraint (Bloem and Bambos 2014).
 In the MDP setting, the above problem is equivalent to finding the maximum likelihood estimate of $\theta$:
 $\hat{\theta}=\operatorname*{arg\,max}_{\theta}\log\prod_{\tau_{j}\in\mathcal{D}}p_{0}(s_{0})\prod_{t=0}^{T}\pi_{\theta}(a_{t}|s_{t})\mathcal{T}(s_{t+1}|s_{t},a_{t}),$ (2)
-with the parameterized policy $\pi_{\theta}$ defined recursively as:
-[…]
-[…]
-[…]
-[…]
-[…]
+with the parameterized policy $\pi_{\theta}$ defined recursively using $Q^{\textit{soft}}$ and $V^{\textit{soft}}$:
+$\pi_{\theta}(a_{t}|s_{t})$ $=\exp(Q^{\textit{soft}}_{\theta}(s_{t},a_{t})-V^{\textit{soft}}_{\theta}(s_{t})),$ (3)
+$Q^{\textit{soft}}_{\theta}$ $=\theta^{\top}f(s_{t},a_{t})+\gamma\operatorname{\mathbb{E}}_{s_{t+1}\sim\mathcal{T}(\cdot|s_{t},a_{t})}[V^{\textit{soft}}_{\theta}(s_{t+1})],$ (4)
+$V^{\textit{soft}}_{\theta}$ $=\log\sum_{a_{t}\in\mathcal{A}}\exp(Q^{\textit{soft}}_{\theta}(s_{t},a_{t})).$ (5)
 The parameters $\theta\in\mathbb{R}^{k}$ specifying the soft Bellman policy, $\pi_{\theta}(a_{t}|s_{t})$, correspond to the dual variables of the feature matching constraint in Eq.
 (1).
-Without loss of generality, if we define the true reward signal as $r(s,a)=\theta^{\star\top}f(s,a)$ for some reward parameter ${\theta}^{\star}\in\mathbb{R}^{k}$ and features $f$, then the soft Bellman policy parameterized by $\theta^{\star}$ achieves the maximum possible return (for more details see (Ziebart, 2010), Theorem 6.2 and Corollary 6.11).
+Without loss of generality, if we define the true reward signal as $r(s,a)=\theta^{\star\top}f(s,a)$ for some reward parameter ${\theta}^{\star}\in\mathbb{R}^{k}$ and features $f$, then the soft Bellman policy parameterized by $\theta^{\star}$ achieves the maximum possible return (for more details see (Ziebart 2010), Theorem 6.2 and Corollary 6.11).
 [^1] To allow an arbitrary reward specification, the IRL objective defined in Eq.
-1 can be expressed in terms of occupancy measures instead of feature expectations (Ho & Ermon, 2016).
+1 can be expressed in terms of occupancy measures instead of feature expectations (Ho and Ermon 2016).
 This formulates an iterative approach to finding the true reward parameter $\theta^{\star}$: given demonstration set $\mathcal{D}$ and your estimate $\theta$, alternate between inferring the policy $\pi_{\theta}(a_{t}|s_{t})$ based on Eq.
@@
 (2).
-This line of work provides many insights that have been used to derive modern IRL (Fu et al., 2018; Reddy et al., 2019; Gao et al., 2018; Garg et al., 2021) and RL (Haarnoja et al., 2017, 2018) algorithms.
+This line of work provides many insights that have been used to derive modern IRL (Fu, Luo, and Levine 2018; Reddy, Dragan, and Levine 2019; Gao et al.
+2018; Garg et al.
+2021) and RL (Haarnoja et al.
+2017, 2018) algorithms.
 Unlike these prior works, we consider the setting where we no longer have access to the expert policy $\pi_{E}$, and instead are given a heterogeneous set of potentially suboptimal demonstrations from which we want to learn a well performing policy.
-To accomplish this, we first describe the model we use to capture the suboptimality of demonstrators in Section 3, following which we frame this problem within the IRL framework and derive our own method IRLEED in Section 4.
+To this end, we first describe the model we use to capture the suboptimality of demonstrators in Section 3, following which we frame this problem within the IRL framework and derive our own method IRLEED in Section 4.
 Finally, we describe our experimental results in Section 5, and conclude our work in Section 6.
 3 Suboptimal Demonstrator Model
-In this work, we consider the case where dataset $\mathcal{D}=\{(i,\mathcal{D}_{i})\}_{i=1}^{N}$ consists of a mixture of $N$ demonstrations with varying quality.
+This work considers the case where dataset $\mathcal{D}=\{(i,\mathcal{D}_{i})\}_{i=1}^{N}$ consists of a mixture of $N$ demonstrations with varying quality.
 Each demonstration $\mathcal{D}_{i}=\{\tau_{j}\}_{j=1}^{M_{i}}$ contains a set of $M_{i}$ trajectories sampled from a distinct, fixed policy $\pi_{i}$.
@@
 To achieve this, it is essential to develop a model for the demonstrators $\pi_{i}$ that accurately reflects their suboptimal behavior relative to the optimal policy $\pi_{E}$.
-A popular framework for modeling decision making is the Boltzmann rationality model, which is widely used in psychology (Baker et al., 2009), economics (Luce, 1959) and RL (Laidlaw & Dragan, 2021).
-In this model, the likelihood of an agent choosing a particular action is proportional to the exponential of the reward associated with that action: $\pi(a|s;r,\beta)\propto\exp(\beta r(s,a))$, where the inverse temperature parameter $\beta$ controls the randomness of the choice.
+A popular framework for modeling decision making is the Boltzmann rationality model, which is widely used in psychology (Baker, Saxe, and Tenenbaum 2009), economics (Luce 1959) and RL (Laidlaw and Dragan 2021).
+In this model, the likelihood of an agent choosing an action is proportional to the exponential of the reward associated with that action: $\pi(a|s;r,\beta)\propto\exp(\beta r(s,a))$, where the inverse temperature parameter $\beta$ controls the randomness of the choice.
 While this formulation provides an intuitive way to capture one form of suboptimality through $\beta$, it traditionally assumes that agents have unbiased knowledge about the true reward $r(s,a)$.
@@
 Within the context of the IRL framework described in Section 2, we recall that the expert policy $\pi_{E}$ aligns with the true reward, formulated as $r(s,a)=\theta^{\star\top}f(s,a)$.
-Applying the aforementioned concept, we can model a demonstrator $i$’s perceived reward as a deviation from the true reward: $r_{i}(s,a)=(\theta^{\star}+\epsilon_{i})^{\top}f(s,a)$, where $\epsilon_{i}\in\mathbb{R}^{k}$.
+Applying the aforementioned concept, we can model demonstrator $i$’s perceived reward as a deviation from the true reward: $r_{i}(s,a)=(\theta^{\star}+\epsilon_{i})^{\top}f(s,a)$, where $\epsilon_{i}\in\mathbb{R}^{k}$.
 By incorporating this altered reward $r_{i}$ into the soft Bellman policy defined in Eq.
 (3), we derive a recursive parameterization for demonstrator $i$’s policy:
-[…]
-[…]
+$\pi_{\theta^{\star},\epsilon_{i},\beta_{i}}(a_{t}|s_{t})\propto\exp(\beta_{i}(Q^{\textit{soft}}_{\theta^{\star}+\epsilon_{i}}(s_{t},a_{t}))),$ (6)
+where $\beta_{i}\in[0,\infty)$, and the normalizing factor not shown above should be modified to account for $\beta_{i}$: $V^{\textit{soft}}_{\theta^{\star}+\epsilon_{i}}(s_{t})=\frac{1}{\beta_{i}}\log\sum_{a_{t}\in\mathcal{A}}\exp(\beta_{i}Q^{\textit{soft}}_{\theta^{\star}+\epsilon_{i}}(s_{t},a_{t}))$.
 Our model captures two distinct aspects of suboptimal behavior: (1) $\epsilon_{i}$ quantifies the demonstrator’s accuracy in estimating the true reward $r$, where $\epsilon_{i}^{\top}f(\tau)$ represents the estimation error.
@@
 Remark 3.1.
-Before moving forward, it is important to note that this model extends outside the scope of problems where the true reward $r$ is an affine transformation of some feature vector $f$.
+It is important to note that this model extends outside the scope of problems where the true reward $r$ is an affine transformation of some feature vector $f$.
 In the general setting, we can parameterize demonstrator $i$’s reward as $r_{i}=r+\epsilon_{i}$, a combination of the true reward $r\in\mathcal{R}$ and their deviation $\epsilon_{i}\in\mathcal{R}$, where $r$ and $\epsilon_{i}$ can be represented by neural networks to allow for general function approximation.
 We opt to formulate our proposed model within the classic feature matching scope, as this facilitates direct comparison with the seminal works used to derive our results in the following section.
-However, our proposed model’s applicability extends beyond this traditional framework, as we elaborate in Section 4.3, and demonstrate through experiments in Section 5.
+However, our proposed model’s applicability extends beyond this traditional framework, as we elaborate in Section 4, and demonstrate through experiments in Section 5.
 Until then, we utilize the feature matching framework described, referencing it as standard IRL.
@@
 Up to this point, we have formulated a model for each demonstrator $i$ based on the parameter $\theta^{\star}$, which defines the true reward, and parameters $\epsilon_{i}$ and $\beta_{i}$, which characterize demonstrator suboptimalities.
-However, we do not have access to any of these parameters, and are instead provided with a dataset of demonstrations $\mathcal{D}$.
+However, we do not have access to these parameters, and are instead provided with a dataset of demonstrations $\mathcal{D}$.
 Recall that standard IRL assumes that all demonstrations come from the optimal policy $\pi_{E}$, and follows the iterative approach described in Section 2 to find it.
-However, applying this approach to the heterogeneous regime will yield an averaged policy, which can lead to subpar performance if a significant portion of the provided demonstrations are suboptimal.
+However, applying this approach to the heterogeneous regime will yield an averaged policy, leading to subpar performance if a significant portion of the demonstrations are suboptimal.
 To refine this, we introduce IRLEED, an innovative extension of the IRL framework tailored for the suboptimal setting.
-Firstly, Section 4.1 details how IRLEED extends the standard IRL framework, which corresponds to the predefined feature matching setting.
-Following this, Section 4.2 presents insightful comparisons between our IRLEED approach and the standard IRL framework, along with the related behavior cloning approach, ILEED.
-Finally, Section 4.3 explains how IRLEED can be applied to the generalized IRL setting, addressing the practical considerations necessary for implementing our framework in conjunction with existing IRL methodologies.
-4.1 IRLEED
-Utilizing the dataset $\mathcal{D}=\{(i,\mathcal{D}_{i})\}_{i=1}^{N}$, we can jointly estimate the unknown parameters by maximizing the likelihood defined below:
+IRLEED
+Utilizing the dataset $\mathcal{D}=\{(i,\mathcal{D}_{i})\}_{i=1}^{N}$, we can jointly estimate the maximum likelihood parameters:
 $\mathcal{L}(\theta,\epsilon,\beta)=\log\prod_{\mathcal{D}_{i}\in\mathcal{D}}\prod_{\tau_{j}\in\mathcal{D}_{i}}P(\tau_{j}|\theta,\epsilon_{i},\beta_{i}),$ (7)
-where $\epsilon=\{\epsilon_{i}\}_{i=1}^{N}$, $\beta=\{\beta_{i}\}_{i=1}^{N}$, and the probability $P(\tau_{j}|\theta,\epsilon_{i},\beta_{i})$ is given by:
-$P(\tau|\theta,\epsilon_{i},\beta_{i})=p_{0}(s_{0})\prod_{t=0}^{T}\pi_{\theta,\epsilon_{i},\beta_{i}}(a_{t}|s_{t})\mathcal{T}(s_{t+1}|s_{t},a_{t}).$ (8)
-Under the specified demonstrator model, this likelihood serves as a suitable loss function. Specifically, rewriting the joint likelihood equivalently only over $\theta$:
-$\mathcal{L}(\theta)=\max_{\epsilon,\beta}\mathcal{L}(\theta,\epsilon,\beta),$ (9)
-it is easy to show that $\mathcal{L}(\theta)$ is a proper objective function.
+where $\epsilon=\{\epsilon_{i}\}_{i=1}^{N}$, $\beta=\{\beta_{i}\}_{i=1}^{N}$, and the probability $P(\tau_{j}|\theta,\epsilon_{i},\beta_{i})$ is given by: $P(\tau|\theta,\epsilon_{i},\beta_{i})=p_{0}(s_{0})\prod_{t=0}^{T}\pi_{\theta,\epsilon_{i},\beta_{i}}(a_{t}|s_{t})\mathcal{T}(s_{t+1}|s_{t},a_{t})$.
+Under the specified demonstrator model, this likelihood serves as a suitable loss function.
+Specifically, rewriting the joint likelihood equivalently only over $\theta$: $\mathcal{L}(\theta)=\max_{\epsilon,\beta}\mathcal{L}(\theta,\epsilon,\beta)$, it is easy to show that $\mathcal{L}(\theta)$ is a proper objective function.
 Proposition 4.1.
 $\theta^{\star}$ is the (non–unique) maximizer of $\mathcal{L}(\theta)$.
-Moreover, this formulation provides us with a convenient way to derive the gradient of likelihood function $\mathcal{L}(\theta,\epsilon,\beta)$.
+This provides us with a convenient way to derive the gradient of likelihood function $\mathcal{L}(\theta,\epsilon,\beta)$.
 Lemma 4.2.
@@
 Proof.
-First, note that parameter $\beta_{i}$ can be folded inside the modified soft Bellman policy of Eq.
+First, the parameter $\beta_{i}$ can be folded into the modified soft Bellman policy of Eq.
 (6) as the reward estimate: $\beta_{i}(\theta+\epsilon_{i})^{\top}f(s_{t},a_{t})$.
 Second, we can express the likelihood as $\mathcal{L}=\sum_{i=1}^{N}\mathcal{L}_{i}$, where $\mathcal{L}_{i}=\log\prod_{\tau_{j}\in\mathcal{D}_{i}}P(\tau_{j}|\theta,\epsilon_{i},\beta_{i})$.
-Setting $u_{i}=\beta_{i}(\theta+\epsilon_{i})$, we know that $\nabla_{u_{i}}\mathcal{L}_{i}=\tilde{f}_{\mathcal{D}_{i}}-\bar{f}_{\pi_{\theta,\epsilon_{i},\beta_{i}}}$ $\forall i\in\{1,\ldots,N\}$ (see (Ziebart, 2010), Theorem 6.2 and Lemma A.2).
-Applying the chain rule completes the proof.
+Setting $u_{i}=\beta_{i}(\theta+\epsilon_{i})$, we know that $\nabla_{u_{i}}\mathcal{L}_{i}=\tilde{f}_{\mathcal{D}_{i}}-\bar{f}_{\pi_{\theta,\epsilon_{i},\beta_{i}}}$ $\forall i\in\{1,\ldots,N\}$ (see (Ziebart 2010), Theorem 6.2, Lemma A.2).
+Using the chain rule completes the proof.
 ∎
-Notably, this result has an intuitive interpretation.
+This result has an intuitive interpretation.
 Recall that for each demonstrator $i$, $\tilde{f}_{\mathcal{D}_{i}}-\bar{f}_{\pi_{\theta,\epsilon_{i},\beta_{i}}}$ represents the difference between the empirical feature vector given samples from their policy, $\tilde{f}_{\mathcal{D}_{i}}$, and the expected feature vector under the probabilistic model, $\bar{f}_{\pi_{\theta,\epsilon_{i},\beta_{i}}}$.
 With this in mind, Eq.
-(11) shows that we update $\epsilon_{i}$ to directly match feature expectation with demonstrator $\pi_{i}$, just like in standard IRL.
+(9) shows that we update $\epsilon_{i}$ to match feature expectation with demonstrator $\pi_{i}$, just like standard IRL.
 On the other hand, Eq.
 (8) shows that we update $\theta$ to match a weighted average over feature expectations provided by all demonstrators, where their respective precision $\beta_{i}$ determines their contribution.
-Finally, Eq.
+Lastly, Eq.
 (10) shows that we update this precision $\beta_{i}$ to balance the expected returns of the demonstrator, $\pi_{i}$, and our probabilistic model, $\pi_{\theta,\epsilon_{i},\beta_{i}}$, under our estimate of the perceived reward, $r_{i}(s,a)=(\theta+\epsilon_{i})^{\top}f(s,a)$.
 This way, when the probabilistic model outperforms the demonstration set under the estimated reward $r_{i}$, we decrease the precision $\beta_{i}$ to lower its relative performance, and vice versa.
-The above formulation gives us a general algorithm that extends the IRL framework described in Section 2.
+This gives us a general algorithm that extends the IRL framework.
 Specifically, IRLEED follows an iterative approach to finding the true reward parameter $\theta^{\star}$: given demonstration set $\mathcal{D}=\{(i,\mathcal{D}_{i})\}_{i=1}^{N}$ and estimates $\theta,\epsilon,\beta$, for each demonstrator $i\in\{1,\ldots,N\}$, alternate between inferring the policy $\pi_{\theta,\epsilon_{i},\beta_{i}}(a_{t}|s_{t})$ based on Eq.
@@
 (8), (9), (10), respectively.
-As aforementioned, many different techniques can be used for both the inference and optimization procedures.
-For example, to infer the policy $\pi_{\theta,\epsilon_{i},\beta_{i}}(a_{t}|s_{t})$ given $\theta,\epsilon_{i},\beta_{i}$, one can use soft versions of the standard value iteration and Q-learning algorithms depending on whether the dynamics $\mathcal{T}$ are provided or not (Bloem & Bambos, 2014).
-To compute feature expectations $\bar{f}_{\pi_{\theta,\epsilon_{i},\beta_{i}}}$ based on the inferred policy, one can directly compute the expectations by utilizing the standard dynamic programming operator when the dynamics $\mathcal{T}$ are known, or estimate the expectations by utilizing Monte Carlo simulations of the MDP when the dynamics $\mathcal{T}$ are unknown.
-4.2 Comparisons to Standard IRL and ILEED
-We can see that this likelihood maximization problem is closely related to the maximum causal entropy IRL framework described in Eqs.
+As aforementioned, many techniques can be used for both the inference and optimization procedures.
+To infer the policy $\pi_{\theta,\epsilon_{i},\beta_{i}}(a_{t}|s_{t})$ given $\theta,\epsilon_{i},\beta_{i}$, one can use soft versions of the standard value iteration and Q-learning algorithms, depending on whether the dynamics $\mathcal{T}$ are provided or not (Bloem and Bambos 2014).
+To compute feature expectations $\bar{f}_{\pi_{\theta,\epsilon_{i},\beta_{i}}}$ based on the inferred policy, one utilize the standard dynamic programming operator when the dynamics $\mathcal{T}$ are known, or estimate the expectations by utilizing Monte Carlo simulations of the MDP when the dynamics $\mathcal{T}$ are unknown.
+Comparisons to Standard IRL and ILEED
+This likelihood maximization problem is closely related to the maximum causal entropy IRL framework described in Eqs.
 (1)- (5).
@@
 (2) It uses a learnable parameter $\beta_{i}$ to tune the magnitude of uncertainty related to demonstrations provided in $\mathcal{D}_{i}$.
-Moreover, IRLEED generalizes the standard IRL framework.
+Moreover, IRLEED generalizes standard IRL.
 Remark 4.3.
-IRLEED recovers the standard IRL framework when we set $\epsilon_{i}=[0]^{k}$ and $\beta_{i}=1$ as constants.
+IRLEED recovers the IRL framework when we set $\epsilon_{i}=[0]^{k}$ and $\beta_{i}=1$ as constants.
 To show the advantage of IRLEED within the suboptimal demonstration setting, we note that the policy recovered by standard IRL can only perform as well as the average demonstration $\mathcal{D}_{i}$ provided.
@@
 Proposition 4.4.
-Given demonstrations $\mathcal{D}=\{(i,\mathcal{D}_{i})\}_{i=1}^{N}$ produced by suboptimal policies $\pi_{i}$ defined according to Eq.
-(6), let $\theta_{IRL}$ denote the naive solution to the standard IRL problem defined by Eq.
+Given $\mathcal{D}=\{(i,\mathcal{D}_{i})\}_{i=1}^{N}$ produced by suboptimal policies $\pi_{i}$ according to Eq.
+(6), let $\theta_{IRL}$ denote the naive solution to the IRL problem defined by Eq.
 2, which treats demonstration set $\mathcal{D}$ as one homogeneous set produced by $\pi_{E}$.
@@
 $\operatorname{\mathbb{E}}_{\pi_{\theta_{IRL}}}[r(s,a)]=\frac{1}{N}\sum_{i=1}^{N}\operatorname{\mathbb{E}}_{\pi_{i}}[r(s,a)]\leq\operatorname{\mathbb{E}}_{\pi_{\theta^{\star}}}[r(s,a)],$ (11)
-where for simplicity, we assume that all demonstrators provide the same number of trajectories $M_{i}=M$.
+where we assume that all demonstrators provide the same number of trajectories $M_{i}=M$.
 As mentioned in Proposition 4.1, IRLEED can recover the true reward parameter $\theta^{\star}$, which outperforms $\theta_{IRL}$ unless the demonstrations provided are optimal.
 Unfortunately, since $\theta^{\star}$ is not the unique solution to IRLEED, we can not make guarantees on recovering the true reward parameter $\theta^{\star}$.
-Nonetheless, our experimental results in Section 5 demonstrate that IRLEED improves performance over IRL frameworks in both the feature matching setting described here, as well as more complex scenarios which include high dimensional input spaces and human demonstrations.
-Finally, we note that when the MDP is deterministic, we can express the policy without recursion:
-$\pi_{\theta,\epsilon_{i},\beta_{i}}(\tau)\propto\exp(\beta_{i}(\theta+\epsilon_{i})^{\top}f(\tau)),$ (14)
-where $f(\tau)=\sum_{(s,a)\in\tau}f(s,a)$, and $\pi(\tau)$ can directly map from trajectories due to the deterministic dynamics $\mathcal{T}$.
-Replacing the probability $P(\tau|\theta,\epsilon_{i},\beta_{i})$ with $\pi_{\theta,\epsilon_{i},\beta_{i}}(\tau)$, we can see that maximizing the likelihood defined by Eq.
-(7) corresponds to the supervised learning approach employed by ILEED (Beliaev et al., 2022).
+Nonetheless, our experimental results in Section 5 demonstrate that IRLEED improves performance over IRL frameworks in both the feature matching setting described here, as well as the general IRL setting, which can be applied to continuous control tasks with high dimensional input spaces.
+Finally, we note that when the MDP is deterministic, we can express the policy without recursion: $\pi_{\theta,\epsilon_{i},\beta_{i}}(\tau)\propto\exp(\beta_{i}(\theta+\epsilon_{i})^{\top}f(\tau)),$ where $f(\tau)=\sum_{(s,a)\in\tau}f(s,a)$, and $\pi(\tau)$ can directly map from trajectories due to the deterministic dynamics $\mathcal{T}$.
+Replacing the probability $P(\tau|\theta,\epsilon_{i},\beta_{i})$ with $\pi_{\theta,\epsilon_{i},\beta_{i}}(\tau)$, maximizing the likelihood defined by Eq.
+(7) corresponds to the supervised learning approach employed by ILEED (Beliaev et al.
+2022).
+Hence unlike ILEED, our framework provides a dynamics–aware solution to the problem of learning from mixtures of suboptimal demonstrations.
 Remark 4.5.
-IRLEED recovers the ILEED framework if we assume all demonstrators $i$ have full knowledge of the true reward, $\epsilon_{i}=[0]^{k}$, and ignore the underlying dynamics of the MDP.
-Hence unlike ILEED, our framework provides us with a dynamics–aware solution to the problem of learning from mixtures of suboptimal demonstrations.
-As we demonstrate in Section 5, IRLEED can be utilized offline to achieve superior performance compared to ILEED.
-4.3 Extension to Generalized IRL
-While the results in the previous sections provide insight on how to learn from suboptimal demonstrations using IRL, relying on hand-crafted features $f$ can hinder our ability to express complex behavior.
+IRLEED recovers the ILEED framework if we assume all demonstrators $i$ have knowledge of the true reward, $\epsilon_{i}=[0]^{k}$, and ignore the dynamics of the MDP.
+Practical Algorithm
+While the results in the previous sections provide insight on how to learn from suboptimal demonstrations using IRL, there are two main concerns: (1) Relying on hand-crafted features $f$ can hinder our ability to express complex behaviors.
+(2) Nothing in the loss prevents the model from overfitting by learning zero shared reward $\theta$, and a completely different reward deviation $\epsilon_{i}$, for each demonstrator $i$.
+To address these concerns, we can extend our approach to generalized IRL:
+$\max_{r\in\mathcal{R}}\min_{\pi\in\Pi}\operatorname{\mathbb{E}}_{\pi_{E}}[r(s,a)]-\operatorname{\mathbb{E}}_{\pi}[r(s,a)]-H(\pi)-\psi(r),$ (12)
+where the reward $r$ belongs to a non restrictive set of functions =$\mathcal{R}=\{r:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}\}$, and $\psi$ is a convex regularizer that is used to prevent overfitting (Ho and Ermon 2016).
+In brief, we extend Eq.
+(12) to the suboptimal setting by using neural networks to parameterize $r$ and $\epsilon=\{\epsilon_{i}\}_{i=1}^{N}$, and implement $\ell_{2}$–regularization on the outputs of $\epsilon$ to mitigate overfitting.
+Our experiment results in Section 5 demonstrate that IRLEED performs well in this generalized setting, and can even be trained offline when paired with Inverse soft-Q learning (IQ), an approach that avoids the iterative process defined by learning a single Q-function, implicitly representing both reward and policy (Garg et al.
+2021).
+We leave the details of this section to Appendix A, covering how IRLEED can be utilized alongside generalized IRL algorithms to address both of the aforementioned concerns.
+Figure 2: We visualize the reward recovered by IRLEED and IRL when trained using suboptimal demonstrations.
+Top left shows the true reward, where the three yellow corners are terminal states.
+Top right shows the normalized state visitation frequency over the entire dataset.
+Bottom left and right show the normalized rewards recovered by IRLEED and IRL respectively.
+We can see that the provided demonstrations are misaligned with the ground truth reward: the state visitation frequency for the top left corner is higher due to demonstrator suboptimalities.
+As expected, the feature matching constraint of IRL absorbs this suboptimality.
+Although the reward recovered by IRL contains information about the true reward, it is incorrectly biasing the top left corner.
+On the other hand, IRLEED is able to remove this bias, providing a better estimate of the ground truth reward.
+Figure 3: We compare the performance of the policies recovered by IRLEED and IRL.
+The left plot shows the relative performance of IRLEED over IRL under varying dataset settings, where the top right corner corresponds to expert data.
+The right plot shows the performance of both policies as we increase the accuracy $\epsilon$ of demonstrators in the dataset, corresponding to the data outlined in red in the left plot.
+On average, IRLEED provided a $30.3\%$ improvement over IRL.
+Setting Cartpole Lunar Lander
+Prec. Acc. Mean IQ IRLEED ILEED IQ IRLEED Mean IQ IRLEED ILEED IQ IRLEED
+L L $0.58$ $0.97$ 1.55 $1.38$ $0.96$ 1.56 $0.46$ $0.68$ 1.28 $0.84$ $0.81$ 1.34
+L H $0.70$ $0.92$ 1.00 $1.00$ $0.96$ 1.00 $0.54$ $0.84$ 1.14 $1.06$ $1.05$ 1.18
+H L $0.64$ $1.03$ 1.09 $1.05$ $0.95$ 1.05 $0.57$ $0.94$ 1.30 $0.85$ $0.99$ 1.14
+H H $1.00$ $0.99$ $1.00$ $0.92$ $1.00$ $1.00$ $0.86$ $0.93$ 0.98 $1.01$ $0.99$ $0.98$
+Table 1: Simulated Control Tasks
+We list the average return of the recovered policies (IQ, IRLEED, ILEED), relative to the best demonstrator’s return, where IQ and IRLEED are used to denote online algorithms.
+This is provided for $4$ dataset settings (Prec., Acc.), where we list the mean performance (Mean) of all demonstrators relative to the best.
+The last row (H,H) corresponds to the clean dataset.
+The return of the best demonstrator for each setting (top to bottom) is $260,500,411,500$ (Cartpole) and $171,213,166,277$ (Lunar Lander).
+5 Experiments
+In this section we evaluate how IRLEED performs when learning from suboptimal demonstrations, using experiments in both online and offline IL settings, with simulated and human-generated data.
+Throughout this section we compared IRLEED to maximum entropy IRL (Ziebart, Bagnell, and Dey 2010), Inverse soft-Q learning (IQ) (Garg et al.
+2021), and ILEED (Beliaev et al.
+2022).
+In total, we performed four sets of experiments: (1) Using a custom Gridworld environment, we simulated suboptimal demonstrations with varying levels of precision and accuracy, and compared the performance and recovered rewards of IRLEED and maximum entropy IRL.
+(2) Using the control tasks, Cartpole, and Lunar Lander, we simulated suboptimal demonstrations with varying levels of expertise and accuracy, and compared the performance of IRLEED, IQ, and ILEED in both the online and offline settings.
+(3) Using the Atari environments, Space Invaders, and Qbert, we combined both simulated and human demonstrations, and compared the performance of IRLEED, IQ, and ILEED in the offline setting.
+(4) Using the continuous control Mujoco task Hopper, we collected suboptimal demonstrations from pretrained policies, and compared the performance of IRLEED and IQ.
+Implementation We utilized the codebases provided by the authors to implement ILEED and IQ, creating IRLEED ontop of the IQ algorithm by adding a learnable parameter $\beta_{i}$, and an additional critic network $\epsilon_{i}$, for each demonstrator $i$.
+Since IQ learns a reward and policy directly by using a single Q function as a critic, we used our additional critic network $\epsilon_{i}$ to directly add bias to the state action function instead of the reward.
+For further implementation details refer to Appendix B.
+1.
+Gridworld For the Gridworld experiments, we implemented IRL and IRLEED by utilizing stochastic value iteration on the reward estimate $\theta$ to infer the policy $\pi$, and Monte Carlo simulations of the MDP to estimate the feature expectations required to optimize $\theta$, repeating both steps until convergence.
+For IRLEED, we updated $\epsilon$ and $\beta$ only after $\theta$ converged, and repeated this for two iterations.
+The reward in this setting was linearly parameterized as $r(s)=\theta^{\top}f(s)$, where $f(s)$ is a one hot encoding for each gridworld state.
+For each level of precision and accuracy, we generated $5$ suboptimal policies according to Equation 6, where $\beta_{i}$ was sampled from a uniform distribution starting at 00, with mean equal to the precision level, and $\epsilon_{i}$ was sampled from a multivariate normal distribution $\mathcal{N}([0]^{k},\mathbf{I}_{k}/\lambda^{2})$, with $\lambda$ equal to the accuracy level.
+We tested $121$ dataset settings, collecting $40$ trajectories from each policy, and using $100$ seeds for each setting.
+Note that we visualized the recovered reward $r(s,a)$ for both IRLEED and IRL under one dataset setting in Figure 2.
+Given demonstrations that are misaligned with the ground truth reward, we saw that the feature matching constraint of IRL absorbed this suboptimality, while IRLEED was able to remove this bias, providing us with a better estimate of the ground truth reward.
+To see the effect of dataset quality on performance, we use Figure 3 to show the relative improvement of the policy recovered by IRLEED over maximum entropy IRL for each dataset setting.
+We can see from the left plot that IRLEED provides improvement over IRL as we decrease accuracy $\epsilon$ (by decreasing $\lambda$), especially when precision $\beta$ is high.
+This signifies the importance of modeling the accuracy $\epsilon$, or reward bias, in addition to the precision $\beta$, or action variance, when learning from suboptimal demonstrations.
+This point is further highlighted in the right plot, which shows that while both methods can achieve good performance under low precision settings, IRLEED can tolerate lower demonstrator accuracy compared to IRL.
+2.
+Simulated Control Tasks For this experiment, each level of precision and accuracy corresponded to $3$ suboptimal policies generated according to $\pi_{i}(a|s)\propto\exp(\beta_{i}(Q(s,a)+\alpha\epsilon_{i}(s,a)))$, where $Q(s,a)$ is a pretrained critic network, $\epsilon_{i}(s,a)$ is a randomly initialized critic network, and $\alpha$ is a scaling factor which is set to zero for the high accuracy setting.
+For the low precision setting, $\beta_{i}$ was sampled from a uniform distribution, whereas for the high precision setting, the stochasticity of the policy was removed by following the maximum state action value.
+We utilized both the online and offline setting, using $30$ seeds for each dataset setting, and displaying our results in Table 1.
+Note that with near binary performance of the recovered policies, the variance is due to demonstration quality, hence we report averages relative to the best demonstrator computed over all initializations.
+We can see that in the online setting, IRLEED outperforms its ablated counterpart, IQ, under all $3$ suboptimal dataset settings (not including the clean dataset), for both environments.
+Furthermore, we see that in the more complex Lunar Lander environment, IQ can not perform as well as the best demonstrator, whereas IRLEED consistently performs on par or better than the best demonstrator.
+To have a fair comparison between IRLEED and its behavior cloning counterpart, ILEED, we repeat these experiments without online access to environment interactions.
+We can see that offline IRLEED outperforms ILEED, under all $3$ suboptimal dataset settings, for both environments.
+Furthermore, although ILEED has comparable performance in the simpler Cartpole environment, it performs poorly in the more complex Lunar Lander environment.
+Moreover, we can see that ILEED shows diminished performance when demonstrators have low accuracy $\epsilon$, signifying that it can not account for the reward biases present in these settings, as was highlighted earlier in Remark 4.5.
+3.
+Atari with Human Data Our demonstrations comprised two unique sources, dataset A: simulated data using a pretrained expert policy and dataset B: collected data using adept human players (Kurin et al.
+2017).
+Since the human dataset B was acquired using screenshots of web-browsers, there is a mismatch in sampling frequencies and colors compared to the simulated environment used in dataset A.
+This provides a challenge akin to crowd-sourcing data from varying sources, where we test if IRLEED can improve IQ when training on the combined dataset AB as opposed to dataset A alone.
+The results shown in Table 2 list the return of the recovered policies for both environments (Space Invaders and Qbert), averaged over $5$ seeds.
+While using dataset A is enough to reach decent performance, all methods do poorly when trained exclusively on dataset B due to the mismatch between the simulation environment and the data.
+However, we can see that IRLEED significantly improves its performance when utilizing the combined dataset AB.
+Overall we see IRLEED shows a $17\%$ improvement over IQ, and a $41\%$ improvement over ILEED in the combined AB setting.
+Dataset Space Invaders Qbert
+ILEED IQ IRLEED ILEED IQ IRLEED
+AB $802$ $755$ 910 $5511$ $8182$ 9336
+A $783$ $754$ $768$ $5926$ $8092$ $8049$
+B $110$ $179$ $123$ 00 $809$ $1087$
+Table 2: Utilizing Demonstrations from Varying Sources
+We list the average return of the recovered policies, using the simulated dataset A, human dataset B (Kurin et al.
+2017), and their combination AB.
+We note that IRLEED maintains similar per-seed variance as IQ on AB - Space Invaders: $61$ vs $54$ (IQ), Qbert: $315$ vs $376$ (IQ).
+Figure 4: We plot the mean episode return of the policies learned by IRLEED (solid) and IQ (dashed) as they train on data from a single source (right) vs a mixture source (left).
+4.
+Mujoco For our results on the Mujoco task Hopper, we collect a single trajectory from three sources, dataset A: a pretrained expert policy (return of $3457$), dataset B: a partially trained policy (return of $2121$) and a barely trained policy C (return of $216$).
+Due to the complex nature of this continuous control tasks, the partially trained policies can not provide adequate demonstrations for learning.
+Hence our goal was to see if IRLEED could improve the capability of IQ when training on the combined dataset ABC by filtering the suboptimal demonstrations.
+The results shown in Figure 4 plot the mean episode return of the policies during training, averaged over $15$ initializations.
+As expected, both methods do poorly when training on dataset C, while achieving improved performance when training on dataset B.
+However, we see IRLEED outperforms IQ when training on the combined dataset ABC since it is more robust to the suboptimal demonstrations present in the data.
+However, we note that there is still a gap between the desired performance we see when training on dataset A alone, which demonstrates the difficulty of the task at hand.
+6 Conclusion
+Summary This paper addresses the challenges of leveraging suboptimal and heterogeneous demonstrations in IL by introducing IRLEED, a novel IRL framework.
+By integrating a general model of demonstrator suboptimality that accounts for reward bias and action variance, with a maximum entropy IRL framework, IRLEED mitigates the limitations inherent in standard IL algorithms, which often fail to account for the diversity and imperfections in real-world data.
+Our key insight is that: supervised learning frameworks should account for the heterogeneity in crowd-sourced datasets by leveraging information about the source identity during training.
+Limitations Although we have shown that IRLEED is a general framework that improves over IRL in a multitude of settings, design choices such as the demonstrator model and reward regularization can be further analyzed.
+While we do try to make our framework easy to interpret, there is inherent complexity added as we need to simultaneously learn multiple demonstrator models.
+We also note that apart from the Gridworld experiments, we did not analyze reward recovery as it is hard to interpret in complex environments.
+However, our work addresses a problem setting where achieving desirable performance is a challenge in itself, and many recovered policies are below optimal.
+Lastly, we have not provided guarantees that state whether IRLEED will recover a better reward estimate compared to IRL.
+While we believe this is a good direction for future work, our goal was to present a novel formulation easily integrated with existing IRL methods, where further theoretical analysis is only possible under additional assumptions, and not crucial given the nonconvexity of the general setting.
+Acknowledgements
+This research was supported by NSF ECCS grant #2419982.
+Appendix A Extension to Generalized IRL
+While the results in Section 4 provide insight on how to learn from suboptimal demonstrations using IRL, relying on hand-crafted features $f$ can hinder our ability to express complex behavior.
 In general, the true reward $r$ can belong to a non restrictive set of functions =$\mathcal{R}=\{r:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}\}$.
@@
 $\max_{r\in\mathcal{R}}\min_{\pi\in\Pi}\operatorname{\mathbb{E}}_{\pi_{E}}[r(s,a)]-\operatorname{\mathbb{E}}_{\pi}[r(s,a)]-H(\pi)-\psi(r),$ (13)
-where $\psi$ is a convex reward regularizer that smoothly penalizes differences between occupancy measures to prevent overfitting (Ho & Ermon, 2016).
+where $\psi$ is a convex reward regularizer that smoothly penalizes differences between occupancy measures to prevent overfitting (Ho and Ermon 2016).
 A naive solution to this nested min-max objective involves an outer loop learning the reward, and an inner loop executing maximum entropy RL with this reward to find the optimal policy.
@@
 Note that under a constant reward regularizer $\psi$, if we express the rewards as $r_{i}(s,a)=(\theta+\epsilon_{i})^{\top}f(s,a)$, then the gradients of the above objective are identical to the ones described in Lemma 4.2.
-In summary, we have shown how any learning algorithm that solves the generalized IRL objective defined in Equation 15 can be utilized alongside our framework by re-paremetrizing the reward and policy accordingly.
-Our experiment results in Section 5 demonstrate that IRLEED performs well in this generalized setting, and can even be trained offline when paired with Inverse soft-Q learning (IQ), an approach that avoids the iterative process defined by learning a single Q-function, implicitly representing both reward and policy (Garg et al., 2021).
-5 Experiments
-In this section, we empirically evaluate how IRLEED performs when learning from suboptimal demonstrations, using experiments in both online and offline IL settings, with simulated and human-generated data.
-In total, we performed three sets of experiments, comparing IRLEED to maximum entropy IRL (Ziebart et al., 2010), Inverse soft-Q learning (IQ) (Garg et al., 2021), and ILEED (Beliaev et al., 2022).
-We go over the results of each experiment below, briefly detailing the corresponding setup for each.
-For further implementation details refer to Appendix A.
-5.1 Gridworld
-For our first experiment, we used a custom gridworld environment, simulating suboptimal demonstrations with varying levels of precision and accuracy.
-Using this environment, we compared the performance of IRLEED against maximum entropy IRL.
-Setup The reward in this setting was linearly parameterized as $r(s)=\theta^{\top}f(s)$, where $f(s)$ is simply a one hot encoding for each gridworld state.
-To implement IRL, we utilized stochastic value iteration on the reward estimate $\theta$ to infer the policy $\pi$, and Monte Carlo simulations of the MDP to estimate the feature expectations required to optimize $\theta$, repeating both steps until convergence.
-We used the same processes to implement IRLEED according to the procedure outlined in Section 4, updating $\epsilon$ and $\beta$ only after $\theta$ converged, and repeating this for two iterations.
-For each level of precision and accuracy, we randomly generated $5$ suboptimal policies according to Equation 6, where $\beta_{i}$ was sampled from a uniform distribution starting at 00, with mean equal to the precision level, and $\epsilon_{i}$ was sampled from a multivariate normal distribution $\mathcal{N}([0]^{k},\mathbf{I}_{k}/\lambda^{2})$, with $\lambda$ equal to the accuracy level.
-We tested a total of $121$ dataset settings, collecting $40$ trajectories from each policy to compose our dataset, and comparing the performance of IRLEED and IRL over $100$ seeds for each setting.
-Results We first compare how IRLEED and IRL recover the reward $r(s,a)$ when provided with suboptimal demonstrations.
-We visualize the averaged result for one dataset setting in Figure 2, which shows the true reward, the dataset quality, and the recovered rewards under both methods.
-We can see that the provided demonstrations are misaligned with the ground truth reward: the state visitation frequency for the top left corner is higher due to demonstrator suboptimalities.
-As expected, the feature matching constraint of IRL absorbs this suboptimality.
-Although the reward recovered by IRL contains information about the true reward, it is now incorrectly biasing the top left corner.
-On the other hand, IRLEED is able to remove this bias, providing us with a better estimate of the ground truth reward.
-To see the effect of dataset quality, we use Figure 3 to show the relative improvement of the policy recovered by IRLEED over maximum entropy IRL for each dataset setting.
-We can see from the left plot that IRLEED provides more improvement over IRL as we decrease accuracy $\epsilon$ (by decreasing $\lambda$), or increase precision $\beta$.
-This signifies the importance of modeling the precision $\epsilon$, or reward bias, in addition to the accuracy $\beta$, or action variance, when learning from suboptimal demonstrations.
-This point is further highlighted in the right plot, which shows that although both IRLEED and IRL suffer when demonstrator accuracy is diminished, IRLEED can tolerate lower demonstrator accuracy compared to IRL.
-In total, IRLEED provided a $30.3\%$ improvement over IRL for the $121$ settings tested.
-Figure 2: We visualize the reward recovered by IRLEED and IRL when trained using suboptimal demonstrations.
-Top left shows the true reward, where the three yellow corners are terminal states.
-Top right shows the normalized state visitation frequency over the entire dataset.
-Bottom left and right show the normalized rewards recovered by IRLEED and IRL respectively.
-Figure 3: We compare the performance of the policies recovered by IRLEED and IRL.
-The left plot shows the relative performance of IRLEED over IRL under varying dataset settings, where the top right corner corresponds to expert data.
-The right plot shows the performance of both policies as we increase the average accuracy $\epsilon$ of demonstrators in the dataset, corresponding to the data outlined in red in the left plot.
-5.2 Simulated Control Tasks
-For our second experiment, we used two control tasks, Cartpole, and Lunar Lander, simulating suboptimal demonstrations with varying levels of expertise and accuracy.
-Using these two environments, we compared the performance of IRLEED, IQ, and ILEED in both the online and offline settings.
-Setup To implement IQ, and setup the training and evaluation procedures, we utilized the codebase provided by the authors.
-We implemented ILEED alongside this codebase, using the training and evaluation procedures.
-For IRLEED, we built on top of the aforementioned IQ implementation by adding a learnable parameter $\beta_{i}$, and an additional critic network $\epsilon_{i}$, for each demonstrator $i$.
-Since IQ learns a reward and policy directly by using a single Q function as a critic, we used our additional critic network $\epsilon_{i}$ to directly add bias to the state action function instead of the reward.
-For this experiment, we used $4$ dataset settings corresponding to the precision and accuracy levels: (high, high), (high, low), (low, high), and (low, low).
-For each setting, we randomly generated $3$ suboptimal policies according to $\pi_{i}(a|s)\propto\exp(\beta_{i}(Q(s,a)+\alpha\epsilon_{i}(s,a)))$, where $Q(s,a)$ is a pretrained critic network, $\epsilon_{i}(s,a)$ is a randomly initialized critic network, and $\alpha$ is a scaling factor which is set to zero for the high accuracy setting.
-For the low precision setting, $\beta_{i}$ was sampled from a uniform distribution, whereas for the high precision setting, the stochasticity of the policy was removed by following the maximum state action value.
-We ran this experiment in both the online and offline setting, comparing the performance of IRLEED, IQ, and ILEED, using $30$ seeds for each dataset setting.
-Results We first compare how IRLEED improves over IQ, with online access to environment interactions.
-We provide these results in Tables 1 and 2, which show the return of the recovered policy relative to the return of the best demonstrator in that dataset.
-We can see that IRLEED outperforms its ablated counterpart, IQ, under all $3$ suboptimal dataset settings (not including the clean dataset), for both environments.
-Furthermore, we see that in the more complex Lunar Lander environment, IQ can not perform as well as the best demonstrator, whereas IRLEED consistently performs on par or better than the best demonstrator.
-To have a fair comparison between IRLEED and its behaviour cloning counterpart, ILEED, we repeat these experiments without online access to environment interactions.
-Firstly, we can see that from Tables 1 and 2 that offline IRLEED outperforms ILEED, under all $3$ suboptimal dataset settings, for both environments.
-We can see that although ILEED has comparable performance in the simpler Cartpole environment, it performs poorly in the more complex Lunar Lander environment.
-Moreover, we can see from Table 2 that ILEED shows diminished performance when demonstrators have low accuracy $\epsilon$, signifying that it can not account for the reward biases present in these settings.
-Table 1: Cartpole
-We list the average return of the recovered policies (IQ, IRLEED, ILEED), relative to the return of the best demonstrator.
-This is provided for $4$ dataset settings (Prec., Acc.), where we list the mean performance of all demonstrators relative to the best demonstrator (Mean).
-The last row (H,H) corresponds to a clean dataset.
-The return of the best demonstrator corresponding to each setting (top to bottom) is: $260,500,411,500$.
-Setting Online Offline Prec.
-Acc.
-Mean IQ IRLEED ILEED IQ IRLEED L L $0.58$ $0.97$ 1.55 $1.38$ $0.96$ 1.56 L H $0.70$ $0.92$ 1.00 $1.00$ $0.96$ 1.00 H L $0.64$ $1.03$ 1.09 $1.05$ $0.95$ 1.05 H H $1.00$ $0.99$ $1.00$ $0.92$ $1.00$ $1.00$
-Table 2: Lunar Lander
-See Table 1.
-The return of the best demonstrator corresponding to each setting (top to bottom) is: $171,213,166,277$.
-Setting Online Offline Prec.
-Acc.
-Mean IQ IRLEED ILEED IQ IRLEED L L $0.46$ $0.68$ 1.28 $0.84$ $0.81$ 1.34 L H $0.54$ $0.84$ 1.14 $1.06$ $1.05$ 1.18 H L $0.57$ $0.94$ 1.30 $0.85$ $0.99$ 1.14 H H $0.86$ $0.93$ 0.98 $1.01$ $0.99$ $0.98$
-5.3 Atari with Human Data
-For our final experiment, we used two Atari environments, Space Invaders, and Qbert, utilizing both simulated and human demonstrations.
-Using these two environments, we compared the performance of IRLEED, IQ, and ILEED in the offline setting.
-Setup Our demonstrations were comprised of two unique sources, dataset A: simulated data using a pretrained expert policy and dataset B: collected data using adept human players (Kurin et al., 2017).
-Since the human dataset B was acquired using screenshots of web-browsers, there is a small mismatch in sampling frequencies and colors when compared to the simulated environment used in dataset A.
-This provides a unique challenge akin to crowd-sourcing data from varying sources, where the goal is to see if we can achieve improved performance when training on the combined dataset AB as opposed to dataset A alone.
-Note that we followed the same steps outlined in the previous Section 5.2 to implement the three methods and setup the training and evaluation procedures.
-The results shown are averaged over $5$ seeds.
-Results With this setup, we would like to test if IRLEED can improve the crowd-sourcing capability of IQ, and compare its performance to ILEED.
-We provide these results for both environments in Table 3, which lists the return of the recovered policies.
-Notably, all methods do poorly when trained exclusively on dataset B, due to the mismatch between the simulation environment and the data.
-On the other hand, utilizing dataset A is enough to reach decent performance.
-However, we can see that IRLEED is the only method that significantly improves its performance when utilizing the combined dataset AB as opposed to dataset A.
-Table 3: Utilizing Demonstrations from Varying Sources
-We list the average return of the recovered policies (IQ, IRLEED, ILEED).
-This is provided for $3$ dataset settings, simulated dataset A, human dataset B (Kurin et al., 2017), and their combination AB.
-Overall we see IRLEED shows a $16\%$ improvement over IQ, and a $41\%$ improvement over ILEED in the combined AB setting.
-Dataset Space Invaders Qbert
-ILEED IQ IRLEED ILEED IQ IRLEED
-AB $802$ $755$ 910 $5511$ $8248$ 9336
-A $783$ $754$ $768$ $5926$ $8092$ $8049$
-B $110$ $179$ $123$ 00 $809$ $1087$
-6 Discussion
-In summary, this paper addresses the challenges of leveraging suboptimal and heterogeneous demonstrations in IL by introducing IRLEED, a novel IRL framework.
-By integrating a general model of demonstrator suboptimality that accounts for reward bias and action variance, with a maximum entropy IRL framework, IRLEED successfully mitigates the limitations inherent in standard IL algorithms, which often fail to account for the diversity and imperfections in real-world data.
-We empirically demonstrate that IRLEED can be used alongside existing IRL techniques to improve their performance in the presence of suboptimal and heterogeneous demonstrations.
-Moreover, we show that IRLEED outperforms its behavior cloning counterpart ILEED in the offline setting, demonstrating the importance of utilizing a dynamics-aware approach that considers reward bias.
-Although we have shown that IRLEED improves over IRL in a multitude of settings, we have not provided guarantees that state whether IRLEED will recover a better reward estimate compared to IRL.
-One direction for future work could be to provide more theoretical analysis comparing IRLEED to IRL under simpler settings with strict assumptions.
-7 Impact Statement
-This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
-Appendix A Implementation Details
+We note that in addition to the reward regularization used in Eq.
+(14), we use $\ell_{2}$-regularization on the output of neural networks $\{\epsilon_{i}\}_{i=1}^{N}$ to model the assumption that each demonstrator $i$ has a reward $r_{i}$ that varies slightly from the common reward $r$.
+One direction for future analysis would be to experiment with other regularization techniques, and see which assumptions are better for different kinds of data.
+Appendix B Implementation Details
 Below we detail the experimental setups utilized in Section 5, and provide information on the computational resources used for our work.
-Note that we provide a working implementation of the Gridworld experiments found in Section 5.1 as part of our supplementary material, and plan to release a public implementation of IRLEED with all experiments and datasets used after publication.
-A.1 Gridworld
-To implement both IRLEED and IRL for the Gridworld experiments in Section 5.1, we utilized stochastic value iteration on the reward estimate $\theta$ to infer the policy $\pi$, and Monte Carlo simulations of the MDP to estimate the feature expectations required to optimize $\theta$, repeating both steps until convergence.
+Note that we provide the implementation of IRLEED used during the Gridworld experiments, while referencing implementations used for other experiments below.
+Gridworld
+To implement both IRLEED and IRL for the Gridworld experiments, we utilized stochastic value iteration on the reward estimate $\theta$ to infer the policy $\pi$, and Monte Carlo simulations of the MDP to estimate the feature expectations required to optimize $\theta$, repeating both steps until convergence.
 As stated, when using IRLEED, we only updated $\epsilon$ and $\beta$ after $\theta$ converged, and repeated this for two iterations.
@@
 We tested a total of $121$ dataset settings, collecting $40$ trajectories from each policy to compose our dataset, and comparing the performance of IRLEED and IRL over $100$ seeds for each setting.
-A.2 Simulated Control Tasks
-To implement IQ and setup the training and evaluation procedures for the simulated control tasks in Section 5.2, we utilized the codebase provided by the authors.
-[^2] https://github.com/Div99/IQ-Learn.
-For all experiments, we utilized the provided hyperparameters for IQ.
-To implement ILEED, we followed the codebase provided by the authors, implementing the modified behaviour cloning loss alongside the IQ implementation.
+Simulated Control Tasks
+To implement IQ and setup the training and evaluation procedures for the simulated control tasks, we utilized the codebase provided by the authors.
+[^2] https://github.com/Div99/IQ-Learn For all experiments, we utilized the provided hyperparameters for IQ.
+To implement ILEED, we followed the codebase provided by the authors, implementing the modified behavior cloning loss alongside the IQ implementation.
 [^3] https://github.com/Stanford-ILIAD/ILEED For IRLEED, we built on top of the aforementioned IQ implementation by adding a learnable parameter $\beta_{i}$, and an additional critic network $\epsilon_{i}$, for each demonstrator $i$.
@@
 For IRLEED and ILEED, $\beta_{i}$ was initialized to the same inverse temperature parameter used in the IQ implementation, and the learning rate was set to $45e-$ 4.
-For IRLEED, the additional critic networks $\epsilon_{i}(s,a)$ were initialized with the same procedure as the critic network $Q(s,a)$, utilized the same learning rate divided by $10$, and utilized the scaling factor $\alpha=0.01$.
+For IRLEED, the additional critic networks $\epsilon_{i}(s,a)$ were initialized with the same procedure as the critic network $Q(s,a)$, utilized the same learning rate divided by $10$, a scaling factor of $\alpha=0.01$, and an $\ell_{2}$-regularization weight of $21e-$ 2.
 We trained all methods using $50,000$ and $100,000$ time steps for Cartpole and Lunar Lander, respectively.
@@
 We ran this experiment in both the online and offline setting, comparing the performance of IRLEED, IQ, and ILEED, using $30$ seeds for each dataset setting.
-A.3 Atari with Human Data
-As aforementioned, the implementation for the Atari experiments in Section 5.3 was identical to the one described above in Section A.2.
+Atari with Human Data
+As aforementioned, the implementation for the Atari experiments was identical to the one described above in Appendix B.
 For training, we utilized $1,000,000$ time steps for all methods.
@@
 For evaluation, we sampled $10$ episodes from the policies recovered during the last $5$ epochs of training, where each epoch corresponded to $5,000$ time steps.
-The two datasets used are publicly available, where dataset A: simulated data using a pretrained expert policy, was provided alongside the original IQ implementation (Garg et al., 2021) and dataset B: collected data using adept human players, was part of a larger human study (Kurin et al., 2017).
+The two datasets used are publicly available, where dataset A: simulated data using a pretrained expert policy, was provided alongside the original IQ implementation (Garg et al.
+2021) and dataset B: collected data using adept human players, was part of a larger human study (Kurin et al.
+2017).
 For Space Invaders, dataset A contained $5$ trajectories with an average return of $1285$, whereas dataset B contained $5$ trajectories with an average return of $1801$.
@@
 We ran this experiments using the three combinations of these datasets, averaging our results over $5$ seeds.
-A.4 Computational Resources
-All of the experiments were performed on an internal cluster containing the 20C/40T Intel Xeon Silver 4114 CPU, 64GB RAM, and 4×GTX-1080 GPUs.
-Obtaining results for the Gridworld experiment with $121$ datasets, each with $100$ seeds, took roughly $3$ days.
-Obtaining results for the simulated control tasks, Cartpole and Lunar Lander, took roughly $4$ days each.
-Training models on the Atari environments took $40$ hours per seed for IQ and IRLEED, and $20$ hours per seed for ILEED.
+Mujoco
+To train on the continuous control Mujoco environments, we used the SAC implementation of IQ provided by the authors.
+For IRLEED, we built on top of the aforementioned IQ implementation by adding a learnable parameter $\beta_{i}$, and appending an additional network $\epsilon_{i}$ to the actor for each demonstrator $i$, keeping the critic network identical.
+All three methods utilized the same architectures, and followed the original hyperparamters used for the IQ implementation.
+For IRLEED, $\beta_{i}$ was initialized to the same inverse temperature parameter used in the IQ implementation, and the learning rate was set to $51e-$ 5.
+The additional actor networks $\epsilon_{i}(s,a)$ were initialized with the same procedure as the actor network $\pi(s,a)$, utilized the same learning rate divided by $10$, a scaling factor of $\alpha=0.01$, and an $\ell_{2}$-regularization weight of $21e-$ 2.
+We trained all methods using $500,000$ time steps.
+For evaluation, we sampled $10$ episodes from the policies recovered during training, abd averaged our result over $15$ initializations.
+To create the different datasets consisting of A, B, and C, we utilized pretrained policies.
+Utilizing the demonstrations provided by the IQ codebase, we ran the IQ algorithm for $1,000,000$ timesteps, $50,000$ timesteps, and $1,000$ timesteps, creating three policies with varying optimality (corresponding to datasets A, B, and C, respectively).
+When creating the combined datasets, each policy was used to generate $1000$ state-action pairs, some containing more than one trajectory.
+Computational Resources
+All of the experiments were performed on an shared cluster containing the 20C/40T Intel Xeon Silver 4114 CPU, 64GB RAM, and 4×GTX-1080 GPUs.